### v3

# Spectral properties of symmetrized AMV operators

Abstract.
The symmetrized Asymptotic Mean Value Laplacian $\tilde{\Delta}$, obtained as limit of approximating operators $\tilde{\Delta}_{r}$, is an extension of the classical Euclidean Laplace operator to the realm of metric measure spaces. We show that, as $r\downarrow 0$, the operators $\tilde{\Delta}_{r}$ eventually admit isolated eigenvalues defined via min-max procedure on any compact uniformly locally doubling metric measure space. Then we prove $L^{2}$ and spectral convergence of $\tilde{\Delta}_{r}$ to the Laplace–Beltrami operator of a compact Riemannian manifold, imposing Neumann conditions when the manifold has a non-empty boundary.
1. Introduction
In the past thirty years, much research has been carried out to extend the classical Euclidean Laplace operator to metric measure spaces: see e.g. [CC00, KMS01, Gig15, AB18]. This paper deals with such an extension, namely the symmetrized Asymptotic Mean Value (AMV) Laplacian, proposed in [MT23], see also [Kok06, MT20, AKS22, AKS23]. The symmetrized AMV Laplacian is set as
$\tilde{\Delta}\coloneqq\lim\limits_{r\downarrow 0}\tilde{\Delta}_{r}$ (1)
where for $\mu$-a.e. $x\in X$,
$\tilde{\Delta}_{r}f(x)\coloneqq\frac{1}{2r^{2}}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)(f(y)-f(x))\mathop{}\!\mathrm{d}\mu(y).$
Here $f$ is a locally integrable function defined on a metric measure space $(X,\mathsf{d},\mu)$. Throughout the paper, $B_{r}(z)$ denotes the metric open ball centered at $z\in X$ with radius $r>0$, the notation $V(z,r)$ stands for $\mu(B_{r}(z))$, and $\fint_{B_{r}(z)}$ is shorthand for $V(z,r)^{-1}\int_{B_{r}(z)}$.
Part of the study on the symmetrized AMV Laplacian consists in finding a relevant meaning to the limit in (1). If this is intended in the $L^{2}$ sense, then the associated spectral convergence can be investigated. This is the goal of the present paper. For any $k\in\mathbb{N}$, set
$\tilde{\lambda}_{k,r}\coloneqq\inf_{V\in\mathcal{G}_{k+1}(L^{2}(X,\mu))}\sup_{f\in V}\frac{\tilde{E}_{r}(f)}{\|f\|_{2}},$ (2)
where $\mathcal{G}_{k+1}(L^{2}(X,\mu))$ is the $(k+1)$-th Grassmannian of $L^{2}(X,\mu)$, and $\tilde{E}_{r}(f)$ is the energy functional naturally associated with $\tilde{\Delta}_{r}$ (Definition 3.10). These form a non-decreasing sequence of non-negative numbers. Our first main result states that these numbers eventually correspond to isolated eigenvalues of $\tilde{\Delta}_{r}$ when $(X,\mathsf{d},\mu)$ is compact and uniformly locally doubling (Definition 2.8).
Theorem 1.
Let $(X,\mathsf{d},\mu)$ be a compact uniformly locally doubling metric measure space. For any integer $k\geq 2$, there exists $r_{k}>0$ such that for any $r\in(0,r_{k})$, the operator $\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues
$0=\lambda_{0}(-\tilde{\Delta}_{r})<\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$
such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$.
Our second main result deals with a smooth manifold $M$ endowed with a smooth Riemannian metric $g$. We write $\Delta_{g}$ for the (negative) Laplace–Beltrami operator of $(M,g)$. We let $m\geq 2$ be the dimension of $M$, and we set
$C_{m}\coloneqq\frac{1}{2}\fint_{\mathbb{B}^{m}_{1}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\xi=\frac{1}{2(m+2)}$ (3)
where $\mathbb{B}^{m}_{1}(0)$ is the unit Euclidean ball of $\mathbb{R}^{m}$. In this context, it follows from the equality between symmetrized and non-symmetrized AMV Laplacian and a simple calculation in normal coordinates that
$\tilde{\Delta}_{r}f(x)\stackrel{{\scriptstyle r\downarrow 0}}{{\to}}C_{m}\Delta_{g}f(x)$ (4)
for any $f\in\mathcal{C}^{2}(M)$ and any interior point $x\in M$, see [MT20, MT23] — the convergence is even locally uniform in the interior of $M$, see [AKS22]. We refer to [MT20, MT23, AKS22, AKS23] for related pointwise results in various settings like Carnot groups or Alexandrov spaces.
In this paper, we are interested in the $L^{2}$ version of (4) with a particular interest in the case where $M$ admits a non-empty boundary $\partial M\neq 0$. In this case, we write $\partial_{\nu}f\in\mathcal{C}^{\infty}(\partial M)$ for the normal derivative of a smooth function $f:M\to\mathbb{R}$, and we define
$\mathcal{C}^{\infty}_{\nu}(M)\coloneqq\{f\in\mathcal{C}^{\infty}(M)\,:\,\partial_{\nu}f=0\}.$ (5)
We see $(M,g)$ as a metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ where $\mathsf{d}_{g}$ and $\mathrm{vol}_{g}$ are the Riemannian distance and volume measure on $M$ associated with $g$. Then our statement reads as follows.
Theorem 2.
Let $(M^{m},g)$ be a compact, connected, smooth Riemannian manifold with a non-empty (resp. empty) boundary $\partial M$. Then for any $f\in\mathcal{C}_{\nu}^{\infty}(M)$ (resp. $\mathcal{C}^{\infty}(M)$), as $r\downarrow 0$,
$\tilde{\Delta}_{r}f\stackrel{{\scriptstyle L^{2}}}{{\longrightarrow}}C_{m}\Delta_{g}f.$
We point out that the boundaryless version of this result is rather easy to obtain, while a non-empty boundary is quite tricky to handle. The Neumann condition in the latter case is crucial to ensure convergence: indeed, the sequence $\tilde{\Delta}_{r}f$ may blow-up if this is not imposed.
After the previous $L^{2}$-convergence result, we address the question of spectral convergence, that is to say, the convergence of the associated eigenvalues and eigenfunctions. In this regard we show that, for any $k\in\mathbb{N}$, the function $r\mapsto\tilde{\lambda}_{k,r}$ is bounded in a neighborhood of 0, as proved in the course of Theorem 1. This ensures that the $k$-th lowest eigenvalue of the operator $\tilde{\Delta}_{r}$, which we denote $\lambda_{k}(-\tilde{\Delta}_{r})$, exists for small enough $r$, and that it coincides with $\tilde{\lambda}_{k,r}$. Let $f_{k,r}$ be an $L^{2}$-normalized eigenfunction of $\tilde{\Delta}_{r}$ associated with $\lambda_{k}(-\tilde{\Delta}_{r})$. Recall that if $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$), a Laplace (resp. Neumann) eigenvalue of $(M,g)$ is a number $\mu\geq 0$ for which there exists an associated eigenfuction $f\in\mathcal{C}^{\infty}(M)$ (resp. $\mathcal{C}_{\nu}^{\infty}(M)$) of $\Delta_{g}$, i.e. $\mu f=\Delta_{g}f$.
Theorem 3.
Let $(M^{m},g)$ be a compact, connected, smooth Riemannian manifold. Assume that $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$). For $k\in\mathbb{N}$, let $\mu_{k}$ be the $k$-th lowest Laplace (resp. Neumann) eigenvalue of $\Delta_{g}$. For any $(r_{n})\subset(0,+\infty)$ such that $r_{n}\to 0$, there exists an $L^{2}$-normalized Laplace (resp. Neumann) eigenfunction $f\in\mathcal{C}^{\infty}(M)$ (resp. $\mathcal{C}_{\nu}^{\infty}(M)$) associated with $\mu_{k}$ such that, up to a subsequence,
$\begin{cases}\lambda_{k}(-\tilde{\Delta}_{r_{n}})\to C_{m}\,\mu_{k},\\ f_{k,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\longrightarrow}}f.\end{cases}$ (6)
We point out that the question of spectral convergence for the Gaussian approximation of the Laplace–Beltrami operator of a compact Euclidean submanifold with boundary was raised in [BQWZ12]. This has been one motivation for the present work: to study this convergence with the intrinsic approximation provided by the symmetrized AMV operators $\tilde{\Delta}_{r}$ instead of the extrinsic Gaussian one.
Acknowledgments.
Both authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. They thank the anonymous reviewers for their numerous and constructive suggestions, and Dmitri Pavlov for clarifying a point related to Lemma 2.1.
2. Averaging-like operators
In this section, we consider a fixed metric measure space, that is to say, a triple $(X,\mathsf{d},\mu)$ where $(X,\mathsf{d})$ is a metric space and $\mu$ is a fully supported regular Borel measure on $(X,\mathsf{d})$ such that
$V(x,r)\coloneqq\mu(B_{r}(x))<+\infty$
for any $x\in X$ and $r>0$, where $B_{r}(x)$ denotes the open ball $\{y\in X:\mathsf{d}(x,y)<r\}$. Notice that for any $x\in X$ and $r>0$,
$V(x,r)>0,$
because $\mu$ is fully supported. Moreover, if $X$ is compact, then
$\mu(X)<+\infty$ (7)
since $\mu$ is finite on any ball of radius the diameter of $X$. We set
$0\leq m(r)\coloneqq\inf_{x\in X}V(x,r)\leq M(r)\coloneqq\sup_{x\in X}V(x,r)\leq+\infty.$
Note that our assumptions yield the following preliminary result.
Lemma 2.1.
$L^{2}(X,\mu)$ is separable.
Proof.
We start by proving that $(X,\mathsf{d})$ is a second countable space. Fix $o\in X$. Given $\epsilon>0$ and $N\in\mathbb{N}$ positive, consider the value given by
$\alpha_{\epsilon,N}=\sup\{\left.\mu\right|_{B_{N}(o)}(\cup_{n}B_{\epsilon}(x_{n})):\{x_{n}\}_{n\in\mathbb{N}}\subset X\},$ (8)
where $\left.\mu\right|_{B_{N}(o)}(\cdot)\coloneqq\mu(\cdot\cap B_{N}(o))$. First we show this supremum is attained. Consider $\delta_{k}\rightarrow 0$ and let $\{x_{n}^{k}\}_{n\in\mathbb{N}}\subset X$ such that
$\left.\mu\right|_{B_{N}(o)}(\cup_{n}B_{\epsilon}(x_{n}^{k}))>\alpha-\delta_{k}.$ (9)
Taking
$\{y_{n}\}_{n\in\mathbb{N}}=\cup_{k}\{x_{n}^{k}\}_{n\in\mathbb{N}}$ (10)
we have that
$\left.\mu\right|_{B_{N}(o)}(\cup_{n}B_{\epsilon}(y_{n}))=\alpha_{\epsilon,N}.$ (11)
[…]
$\mu(K)>0.$ (12)
Since we can cover $K$ by a finite number of balls $B_{\epsilon}(z_{k})$, there must exist some $z=z_{k}$ such that
$\mu(B_{\epsilon}(z)\cap K)>0.$ (13)
We then have
$\left.\mu\right|_{B_{N}(o)}\left(\cup_{n}B_{\epsilon}(y_{n})\cup B_{\epsilon}(z)\right)$ $\geq\left.\mu\right|_{B_{N}(o)}\left(\cup_{n}B_{\epsilon}(y_{n})\cup\left(B_{\epsilon}(z)\cap K\right)\right)$ (14)
$=\mu\left(\cup_{n}B_{\epsilon}(y_{n})\right)+\mu\left(B_{\epsilon}(z)\cap K\right)>\alpha_{N,\epsilon}.$ (15)
And so $\{y_{n}\}_{n\in\mathbb{N}}\cup\{z\}$ contradicts the maximality of $\{y_{n}\}_{n\in\mathbb{N}}$. This shows that $\alpha_{\epsilon,N}=\mu(B_{N}(o))$. Since $\cup_{n}B_{\epsilon}(y_{n})\cap B_{N}(o)$ has full measure in the support $B_{N}(o)$ of $\mu|_{B_{N}(o)}$, it is a dense subset of $B_{N}(o)$. This implies that $\cup_{n}B_{2\epsilon}(y_{n})$ is a countable cover of $B_{N}(o)$. To build a countable basis of $X$, consider a sequence $\delta_{k}\rightarrow 0$. For any $k$, take $\cup_{n\in\mathbb{N}}B_{\delta_{k}}(y_{n}^{k,N})$ a countable cover of $B_{N}(z)$. Then the set given by
$\mathcal{B}=\cup_{k,N\in\mathbb{N}}\{B_{\delta_{k}}(y_{n}^{k,N})\}_{n\in\mathbb{N}}$ (16)
is a countable basis of $X$. Given this basis $\mathcal{B}$ we can create a new basis given by the finite union of elements of $\mathcal{B}$, and we call this new basis $\mathcal{B}^{\prime}$ which will also be countable. In particular we have that given some open set $V\subset X$ we can find a sequence $V_{n}$ such that
$V_{n}\subset V_{n+1},\quad\quad\cup_{n}V_{n}=V.$ (17)
This is the case since $\mathcal{B}$ is a countable basis, we can find elements $B_{k}\in\mathcal{B}$ such that
$\cup_{k}B_{k}=V.$ (18)
We conclude by taking $V_{n}=\cup_{k=1}^{n}B_{k}\in\mathcal{B}^{\prime}$. To construct our dense subset of $L^{2}(X,\mu)$ we take finite sums with rational coefficients of the characteristic functions $\chi_{V}$, with $V\in\mathcal{B}^{\prime}$. To show that this is dense in $L^{2}(X,\mu)$ we only need to show that we can approximate arbitrarily well simple functions $\chi_{U}$ where $U\subset X$ is open and $\mu(U)<\infty$ since the measure $\mu$ is outer regular. Given such an open set $U$, take $\chi_{U_{n}}$ where $U_{n}\in\mathcal{B}^{\prime}$ and $U_{n}\subset U_{n+1}$ and $\cup_{n}U_{n}=U$. Then we have by dominated convergence
$\chi_{U_{n}}\rightarrow_{L^{2}(X,\mu)}\chi_{U},$ (19)
concluding the proof. ∎
2.1. Averaging operator
For any $x,y\in X$ and $r>0$, set
$a_{r}(x,y)\coloneqq\frac{1_{B_{r}(x)}(y)}{V(x,r)}\,\cdot$
Consider $u\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $u$ is $\mu$-integrable on $B_{r}(x)$, set
$A_{r}u(x)\coloneqq\fint_{B_{r}(x)}u\mathop{}\!\mathrm{d}\mu=\int_{X}a_{r}(x,y)u(y)\mathop{}\!\mathrm{d}\mu(y).$
Notice that, since $u$ is locally integrable, for any $x\in X$ there exists $r_{x}>0$ such that $A_{r_{x}}u(x)$ is well-defined. However, there may be no uniform $r>0$ for which the integral $A_{r}u(x)$ is well-defined for every $x\in X$.
Let us also set
$a^{*}_{r}(x,y)\coloneqq a_{r}(y,x)=\frac{1_{B_{r}(x)}(y)}{V(y,r)}$
for any $x,y\in X$ and $r>0$. Consider $u\in L^{0}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $v$ is $\mu$-integrable on $B_{r}(x)$, set
$A^{*}_{r}u(x)\coloneqq\int_{B_{r}(x)}\frac{u(y)\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}=\int_{X}a_{r}^{*}(x,y)u(y)\mathop{}\!\mathrm{d}\mu(y).$
Notice that, just like $A_{r}u(x)$, $A^{*}_{r}u(x)$ may not make sense uniformly with respect to $x\in X$.
For any $r>0$, we introduce the following conditions:
$\|A_{r}^{*}1\|_{\infty}<+\infty,$ ($\mathrm{I}_{r}$)
$V(\cdot,r)^{-1}\in L^{1}(X,\mu).$ ($\mathrm{II}_{r}$)
Note that ($\mathrm{II}_{r}$) implies ($\mathrm{I}_{r}$) since
$\|A_{r}^{*}1\|_{\infty}=\sup_{x\in X}|A_{r}^{*}1(x)|=\sup_{x\in X}\int_{B_{r}(x)}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\leq\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\,\cdot$
In the next lemma, we discuss the boundedness and the compactness of the averaging operator $A_{r}$ acting on Lebesgue spaces.
Lemma 2.2.
Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then for any $p\in[1,+\infty]$ the linear operator $A_{r}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{1/p}\cdot$
Moreover, if ($\mathrm{II}_{r}$) holds, then $A_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
Proof.
The case $p=+\infty$ is obvious and holds regardless of ($\mathrm{I}_{r}$). Let us assume that $p<+\infty$. Let $u\in L^{p}(X,\mu)$. By Jensen’s inequality, for any $x\in X$,
$|A_{r}u(x)|^{p}\leq\left(\fint_{B_{r}(x)}|u|^{p}\mathop{}\!\mathrm{d}\mu\right).$
Thus
$\|A_{r}u\|_{p}^{p}$ $\leq\int_{X}\frac{1}{V(x,r)}\int_{B_{r}(x)}|u(y)|^{p}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}\int_{X}\frac{1}{V(x,r)}\underbrace{1_{B_{r}(x)}(y)}_{=1_{B_{r}(y)}(x)}|u(y)|^{p}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}|u(y)|^{p}\underbrace{\int_{B_{r}(y)}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}}_{=A_{r}^{*}1(y)}\mathop{}\!\mathrm{d}\mu(y)\leq\|A_{r}^{*}1\|_{\infty}\|u\|_{p}^{p}$
where we have used the Fubini–Tonelli theorem to get the second equality and ($\mathrm{I}_{r}$) for the last inequality.
Let us now assume that ($\mathrm{II}_{r}$) holds. Since
$\int_{X}\int_{X}a_{r}^{2}(x,y)\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)=\int_{X}\frac{1}{V(x,r)}\fint_{B_{r}(x)}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)=\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}$
we obtain that $A_{r}$ is a Hilbert-Schmidt integral operator acting on the separable space $L^{2}(X,\mu)$ (recall Lemma 2.1); in particular, $A_{r}$ is compact [RS80, Section IV.6]. ∎
In the next statement, we provide an alternative way to prove the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to itself. This goes through the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to the space of continuous functions $\mathcal{C}(X)$ which we obtain for compact spaces $X$ satisfying the following condition:
$\sup_{x\in X}\mu(S_{r}(x))=0$ ($\mathrm{S}_{r}$)
where $S_{r}(x)\coloneqq\{y\in X:\mathsf{d}(x,y)=r\}$.
Lemma 2.3.
Assume that $(X,\mathsf{d},\mu)$ is compact and satisfies ($\mathrm{S}_{r}$) for some $r>0$. Then $A_{r}:L^{2}(X,\mu)\to\mathcal{C}(X)$ is compact and satisfies
$\|A_{r}\|_{2\to\infty}\leq\frac{1}{m(r)^{1/2}}\,\cdot$ (20)
Proof.
We start by noticing that if $u\in L^{2}(X,\mu)$, then $A_{r}(u)$ is continuous. This follows from $V(\cdot,r)^{-1}$ and $\int_{B_{r}(x)}u(y)d\mu(y)$ being continuous. The former holds by assumption. To prove the latter, assume that $\|u\|_{L^{2}(X)}=1$. Then for any $x,z\in X$,
$\left|\int_{B^{r}(x)}u(y)d\mu(y)-\int_{B^{r}(z)}u(y)d\mu(y)\right|$ $\leq\left\|1_{B_{r}(x)}-1_{B_{r}(z)}\right\|_{L^{2}(X)}\left\|u\right\|_{L^{2}(X)}$ (21)
$\leq\mu(B_{r+d(x,z)}(x)-B_{r-d(x,z)}(x))^{\frac{1}{2}},$ (22)
and $\mu(B_{r+d(x,z)}(x)-B_{r-d(x,z)}(x))\to 0$ as $\mathsf{d}(x,z)\to 0$ due to ($\mathrm{S}_{r}$). Moreover, the bound (20) is obtained via Hölder’s inequality: for any $x\in X$,
$|A_{r}u(x)|\leq\left(\fint_{B_{r}(x)}u^{2}\mathop{}\!\mathrm{d}\mu\right)^{1/2}\leq\frac{1}{m(r)^{1/2}}\,\cdot$
To prove compactness, consider $\{f_{n}\}\subset L^{2}(X,\mu)$ such that $\sup_{n}\|f_{n}\|_{2}\leq 1$. Uniform boundedness of $\{A_{r}(f_{n})\}$ follows from (20), and equicontinuity can be obtained by using the inequality (22) applied to the sequence. By the Ascoli-Arzelà Theorem, we can extract from $\{A_{r}(f_{n})\}$ a subsequence which converges in $\mathcal{C}(X)$, concluding the proof. ∎
2.2. Adjoint
Let us focus now on the boundedness and the compactness of the adjoint operator $A_{r}^{*}$. We begin with a simple observation.
Lemma 2.4.
The operator $A_{r}^{*}:L^{1}(X,\mu)\to L^{1}(X,\mu)$ is a contraction for any $r>0$.
Proof.
For any $u\in L^{1}(X,\mu)$,
$\int_{X}|A_{r}^{*}u(x)|\mathop{}\!\mathrm{d}\mu(x)$ $\leq\int_{X}\int_{B_{r}(x)}\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}\int_{X}1_{B_{r}(x)}(y)\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}\left(\int_{X}1_{B_{r}(y)}(x)\mathop{}\!\mathrm{d}\mu(x)\right)\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)=\int_{X}|u(y)|\mathop{}\!\mathrm{d}\mu(y),$
where we used the Fubini–Tonelli theorem to get the penultimate equality. ∎
We continue with the next lemma which covers the case $p>1$.
Lemma 2.5.
Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then for any $p\in[1,+\infty]$, the linear operator $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}^{*}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{(p-1)/p}$
Moreover, this operator is the adjoint of $A_{r}:L^{q}(X,\mu)\to L^{q}(X,\mu)$ for $q\in[1,+\infty]$ such that $1/p+1/q=1$. Lastly, if ($\mathrm{II}_{r}$) holds, then the operator $A_{r}^{*}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
Proof.
For the proof of the first assertion, consider $u\in L^{\infty}(X,\mu)$. Thanks to ($\mathrm{I}_{r}$), for $\mu$-a.e. $x\in X$,
$|A_{r}^{*}u|(x)$ $\leq\int_{B_{r}(x)}\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq\|u\|_{\infty}\int_{B_{r}(x)}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\leq\|A_{r}^{*}1\|_{\infty}\|u\|_{\infty}.$
Thus $A_{r}^{*}:L^{\infty}(X,\mu)\to L^{\infty}(X,\mu)$ is bounded with $\|A_{r}^{*}\|_{\infty\to\infty}\leq\|A_{r}^{*}1\|_{\infty}$. The conclusion for $p\in(1,+\infty)$ follows from the Riesz-Thorin theorem and Lemma 2.4.
Let us prove that $A_{r}$ and $A_{r}^{*}$ are adjoint of each other. Consider $u\in L^{p}(X,\mu)$ and $v\in L^{q}(X,\mu)$. Then
$\int_{X}A_{r}^{*}u(x)\,v(x)\mathop{}\!\mathrm{d}\mu(x)$ $=\int_{X}\int_{X}1_{B_{r}(x)}(y)\,\frac{u(y)}{V(y,r)}\,v(x)\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}\frac{u(y)}{V(y,r)}\int_{X}1_{B_{r}(x)}(y)\,\,v(x)\mathop{}\!\mathrm{d}\mu(x)\mathop{}\!\mathrm{d}\mu(y)$
$=\int_{X}u(y)A_{r}v(y)\mathop{}\!\mathrm{d}\mu(y)$
where we used the Fubini–Tonelli to get the second equality, and the equality $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ to get the last one. As for the compactness of $A_{r}^{*}$ under ($\mathrm{II}_{r}$), this result is a direct consequence of the Schauder theorem for compact operators which can be applied thanks to Lemma 2.2. ∎
2.3. Discussion on the assumptions
Let us discuss the validity of ($\mathrm{I}_{r}$) and ($\mathrm{II}_{r}$). Recall first that : ($\mathrm{II}_{r}$) $\Rightarrow$ ($\mathrm{I}_{r}$). Both properties hold on totally bounded spaces, as seen in the next lemma.
Lemma 2.6.
Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{II}_{r}$) (and then ($\mathrm{I}_{r}$)) holds for any $r>0$.
Proof.
Consider $r>0$ and a finite cover $\{B_{r/2}(x_{i})\}$ of $X$. For any $x\in X$, there exists $i$ such that $x\in B_{r/2}(x_{i})$. Then $B_{r}(x)$ contains $B_{r/2}(x_{i})$ so $V(x,r)\geq V(x_{i},r/2)\geq\min_{j}V(x_{j},r/2)>0$. Thus
$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{\min_{j}V(x_{j},r/2)}<+\infty.$
Remark 2.7.
If $\mu(X)=+\infty$ and $M(r)<+\infty$ then ($\mathrm{II}_{r}$) cannot hold :
$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\geq\frac{\mu(X)}{M(r)}=+\infty.$
This happens, for instance, on $\mathbb{R}^{n}$ endowed with the Euclidean distance and the Lebesgue measure. More generally, this property cannot hold on a locally compact topological group endowed with a left-invariant metric compatible with the Haar measure and with infinite volume (see [Str74, Lemma 1] for more details about these spaces).
If $\mu(X)<\infty$ and $m(r)>0$, then ($\mathrm{II}_{r}$) always holds :
$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{m(r)}<+\infty.$
In this regard, observe that if $X$ is not totally bounded, then there exist $r>0$ small enough and a countable family of disjoint balls $\{B_{r}(x_{i})\}$ in $X$, so that $\mu(X)<\infty$ and $m(r)>0$ cannot hold simultaneously:
$\mu(X)\geq\sum_{i}V(x_{i},r).$
Let us now focus on ($\mathrm{I}_{r}$). We show below that this condition holds on so-called doubling spaces. Let us recall this classical property and its uniform local variant, see e.g. [HKST15] for more details.
Definition 2.8.
The space $(X,\mathsf{d},\mu)$ is called globally doubling if there exists $C>0$ such that for any $x\in X$ and $r>0$,
$V(x,2r)\leq CV(x,r).$ (23)
It is called uniformly locally doubling if there exist $C,r_{0}>0$ such that (23) holds for any $x\in X$ and $r\in(0,r_{0})$.
The celebrated Bishop–Gromov theorem (see e.g. [Cha95, Theorem III.4.5]) implies that any complete Riemannian manifold with a uniform lower bound on the Ricci curvature is uniformly uniformly locally doubling, and globally doubling if the uniform bound is non-negative. This is also true for metric spaces with generalized sectional curvature bounded from below in the sense of Alexandrov [BBI01, Theorem 10.6.6] and $\mathrm{CD}(K,N)$ metric measure space [Vil09, Corollary 30.14].
The next lemma relates the uniformly local doubling condition with ($\mathrm{I}_{r}$).
Lemma 2.9.
Let $(X,\mathsf{d},\mu)$ be uniformly locally doubling with parameters $C,r_{0}$. Then ($\mathrm{I}_{r}$) holds with $\|A_{r}^{*}1\|_{\infty}\leq C$ for any $r\in(0,r_{0})$.
Proof.
For any $x\in X$ and $r\in(0,r_{0})$, the triangle inequality yields that $B_{r}(x)\subset B_{2r}(y)$ for any $y\in B_{r}(x)$. Then
$A_{r}^{*}1(x)=\fint_{B_{r}(x)}\frac{V(x,r)}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq\fint_{B_{r}(x)}\frac{V(y,2r)}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq C.$
Remark 2.10.
Of course if $(X,\mathsf{d},\mu)$ is globally doubling with constant $C$, then ($\mathrm{I}_{r}$) holds with $\|A_{r}^{*}1\|_{\infty}\leq C$ for any $r>0$.
Remark 2.11.
The previous result notably implies that ($\mathrm{I}_{r}$) may hold in situations where ($\mathrm{II}_{r}$) does not. This happens e.g. on a non-compact Riemannian manifold $(M,g)$ with non-negative Ricci curvature endowed with its canonical Riemannian distance $\mathsf{d}$ and volume measure $\mu$. Indeed, such a space has infinite volume, and the Bishop–Gromov theorem implies that $M(r)\leq\mathbb{V}^{n}(r)$ for any $r>0$, where $\mathbb{V}^{n}(r)$ is the Lebesgue measure of an Euclidean ball of radius $r$ in $\mathbb{R}^{n}$. From Remark 2.7, we get that $M$ cannot satisfy ($\mathrm{II}_{r}$) for any $r>0$, while it does satisfies ($\mathrm{I}_{r}$) thanks to the global doubling condition.
We conclude this discussion with two final remarks. First, if $m(r)>0$ and $M(r)<+\infty$, then ($\mathrm{I}_{r}$) always holds with
$\|A_{r}^{*}1\|_{\infty}\leq\frac{M(r)}{m(r)}\,\cdot$
This happens on locally compact topological groups endowed with a left-invariant distance and their Haar measure, compare with Remark 2.7. Secondly, ($\mathrm{I}_{r}$) can easily be seen as a weak variant of the comparability conditions introduced in [Ald18, MT23].
2.4. Symmetrization
For any $x,y\in X$ and $r>0$, set
$\tilde{a}_{r}(x,y)$ $=\frac{1}{2}\left(a_{r}(x,y)+a^{*}_{r}(x,y)\right)$
$=\frac{1}{2}\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)1_{B_{r}(x)}(y)$
Consider $u\in L^{1}_{\text{loc}}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $u$ and $v$ are $\mu$-integrable on $B_{r}(x)$, set
$\tilde{A}_{r}u(x)\coloneqq\frac{1}{2}(A_{r}u(x)+A_{r}^{*}u(x))=\int_{X}\tilde{a}_{r}(x,y)u(y)\mathop{}\!\mathrm{d}\mu(y).$ (24)
Then the next lemma is an obvious consequence of Lemma 2.2 and Lemma 2.5.
Corollary 2.12.
Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is a self-adjoint operator such that
$\|\tilde{A}_{r}\|_{2\to 2}\leq\|A_{r}^{*}1\|_{\infty}^{1/2}.$
Moreover, if ($\mathrm{II}_{r}$) holds, then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
3. Symmetrized AMV operators
In this section, we provide our working definition of the symmetrized AMV $r$-Laplace operator $\tilde{\Delta}_{r}$ and we derive several spectral properties in a general setting.
3.1. Definitions
For this subsection, we consider a metric measure space $(X,\mathsf{d},\mu)$ satisfying ($\mathrm{I}_{r}$) for some fixed $r>0$.
Definition 3.1.
The symmetrized AMV $r$-Laplace operator of $(X,\mathsf{d},\mu)$ is
$\tilde{\Delta}_{r}\coloneqq\frac{1}{r^{2}}(\tilde{A}_{r}-[\tilde{A}_{r}1]\mathrm{I})$
where we recall that $\tilde{A}_{r}$ is defined in (24).
Remark 3.2.
We may use the notation $\tilde{\Delta}_{r,\mathfrak{X}}$ to specify that we work on the metric measure space $\mathfrak{X}=(X,\mathsf{d},\mu)$.
Lemma 3.3.
$\tilde{\Delta}_{r}$ is a bounded, self-adjoint operator acting on $L^{2}(X,\mu)$ with
$\|\tilde{\Delta}_{r}\|_{2\to 2}\leq\frac{1}{2r^{2}}\left(2\|A_{r}^{*}1\|_{\infty}^{1/2}+\|A_{r}^{*}1\|_{\infty}+1\right).$ (25)
Proof.
The self-adjointness of $\tilde{\Delta}_{r}$ is obvious because $\tilde{A}_{r}$ and $\tilde{A}_{r}1]\mathrm{I}$ are self-adjoint too. The boundedness is a consequence of Lemma 2.12. Indeed,
$\|\tilde{\Delta}_{r}\|_{2\to 2}$ $\leq\frac{1}{r^{2}}\left(\|\tilde{A}_{r}\|_{2\to 2}+\|[\tilde{A}_{r}1]I\|_{2\to 2}\right)$
$\leq\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\|\tilde{A}_{r}1\|_{\infty}\underbrace{\|I\|_{2\to 2}}_{=1}\right)$
$\leq\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\frac{\|A_{r}1\|_{\infty}+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
$=\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\frac{1+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
hence $\tilde{\Delta}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is bounded and (25) holds. ∎
Definition 3.4.
The energy functional $\tilde{E}_{r}$ of $(X,\mathsf{d},\mu)$ is the quadratic form on $L^{2}(X,\mu)$ defined by
$\tilde{E}_{r}(f)\coloneqq\frac{1}{4}\int_{X}\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\left(\frac{f(x)-f(y)}{r}\right)^{2}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
The associated bilinear form, which we still denote by $\tilde{E}_{r}$, is given by
$\tilde{E}_{r}(f,\psi)$ $\coloneqq\frac{1}{4}\int_{X}\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)$
$\qquad\qquad\qquad\qquad\times\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
Remark 3.5.
A suitable use of the Fubini–Tonelli theorem shows that the energy functional $\tilde{E}_{r}(f)$ equals the approximate Korevaar–Schoen energy [KS93]
$\frac{1}{2}\int_{X}\fint_{B_{r}(x)}\frac{|f(y)-f(x)|^{2}}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
Remark 3.6.
We may also use the notation $\tilde{E}_{r,\mathfrak{X}}$ to specify the metric measure space $\mathfrak{X}=(X,\mathsf{d},\mu)$.
The next lemma goes back to [AKS23, Lemma 3.1]. We provide a quick proof for completeness.
Lemma 3.7.
For any $f,\psi\in L^{2}(X,\mu)$,
$\tilde{E}_{r}(f,\psi)=\langle-\tilde{\Delta}_{r}f,\psi\rangle_{L^{2}}.$ (26)
Proof.
Note that
$\tilde{E}_{r}(f,\psi)$ $=\frac{1}{4}\int_{X}\int_{B_{r}(x)}\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\frac{(f(x)-f(y))\psi(x)}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$-\frac{1}{4}\int_{X}\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\frac{(f(x)-f(y))\psi(y)}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
Using $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ and then the Fubini theorem, we can rewrite the second term as the opposite of the first one, so that we eventually get (26). ∎
Remark 3.8.
Observe that (26) implies that $\tilde{\Delta}_{r}$ is a non-negative operator, since for any $f\in L^{2}(X,\mu)$,
$\langle-\tilde{\Delta}_{r}f,f\rangle_{L^{2}}=\tilde{E}_{r}(f)\geq 0$
Let us recall the definition of spectrum.
Definition 3.9.
We let $\sigma(-\tilde{\Delta}_{r})$ denote the spectrum of $\tilde{\Delta}_{r}$, that is to say, the set of elements $\lambda\in\mathbb{C}^{*}$ such that $\tilde{\Delta}_{r}-\lambda\mathrm{I}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is not a bijection.
It is well-known from classical functional analysis that the spectrum $\sigma(T)$ of a bounded operator $T$ acting on a Banach space $E$ can be decomposed as
$\sigma(T)=\sigma_{\text{p}}(T)\cup\sigma_{\text{c}}(T)\cup\sigma_{\text{a}}(T)$
where:
  • $\sigma_{\text{p}}(T)$ is the point spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ such that $(T-\lambda\mathrm{I})f=0$ for some non-zero $f\in E$, in which case $\lambda$ is called an eigenvalue and $f$ an eigenvector of $T$,
  • $\sigma_{\text{c}}(T)$ is the compression spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ whose conjugate $\bar{\lambda}$ is an eigenvalue of the adjoint $T^{*}$,
  • $\sigma_{\text{a}}(T)$ is the approximate point spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ for which there exists $(f_{n})\subset E$ with $\|f_{n}\|=1$ for any $n$ such that $\|(T-\lambda\mathrm{I})f_{n}\|\to 0$.
Since $\tilde{\Delta}_{r}$ is self-adjoint and non-negative, we know that
$\sigma(-\tilde{\Delta}_{r})\subset[0,+\infty].$
This implies that $\sigma_{\text{p}}(-\tilde{\Delta}_{r})=\sigma_{\text{c}}(-\tilde{\Delta}_{r})$, so that
$\sigma(-\tilde{\Delta}_{r})=\sigma_{\text{p}}(-\tilde{\Delta}_{r})\cup\sigma_{\text{a}}(-\tilde{\Delta}_{r}).$ (27)
Definition 3.10.
For any $k\in\mathbb{N}$, we define
$\tilde{\lambda}_{k,r}\coloneqq\inf_{V\in\mathcal{G}_{k+1}(L^{2}(X,\mu))}\sup_{f\in V}\frac{\tilde{E}_{r}(f)}{\|f\|_{2}},$ (28)
where $\mathcal{G}_{k+1}(L^{2}(X,\mu))$ is the $(k+1)$-th Grassmannian of $L^{2}(X,\mu)$.
Remark 3.11.
Let $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ denote the essential spectrum of $\tilde{\Delta}_{r}$, i.e. the closed subset of $\sigma(-\tilde{\Delta}_{r})$ made of those $\lambda$ such that $\tilde{\Delta}_{r}-\lambda\mathrm{I}$ is not a Fredholm operator. Since $\tilde{\Delta}_{r}$ is self-adjoint, the Fischer–Polyà minimum-maximum principle (see e.g. [WS72, p.12]) asserts that if there exists a positive integer $N$ such that $\tilde{\lambda}_{N,r}<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ then $\tilde{\Delta}_{r}$ admits $N+1$ isolated eigenvalues
$\lambda_{0}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{N}(-\tilde{\Delta}_{r})<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$
such that for any $k\in\{0,\ldots,N\}$,
$\tilde{\lambda}_{k,r}=\lambda_{k}(-\tilde{\Delta}_{r}).$
3.2. Spectral properties
Our first spectral result on $\tilde{\Delta}_{r}$ is the following. Note that we need the compactness of $\tilde{A}_{r}$ here, thus we assume ($\mathrm{II}_{r}$).
Proposition 3.12.
Let $(X,\mathsf{d},\mu)$ be a metric measure space satisfying ($\mathrm{II}_{r}$) for some fixed $r>0$. Assume that $\lambda\in\sigma(-\tilde{\Delta}_{r})$ satisfies
$\lambda<\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}}.$ (29)
Then $\lambda$ is an isolated eigenvalue of $\tilde{\Delta}_{r}$ which does not belong to $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$.
Proof.
Let us first show that $\lambda$ is an eigenvalue, that is to say, that $\lambda$ belongs to the point spectrum. According to (27), it is enough to show that if $\lambda$ is in the approximate point spectrum, then it is in the point spectrum. If this is the case, then there exists a sequence $(f_{n})\in L^{2}(X,\mu)$ such that $\|f_{n}\|_{L^{2}(X)}=1$ and
$\|-\tilde{\Delta}_{r}f_{n}-\lambda f_{n}\|_{L^{2}(X)}=\left\|\left(\frac{[\tilde{A}_{r}1]}{r^{2}}-\lambda\right)f_{n}-\frac{\tilde{A}_{r}f_{n}}{r^{2}}\right\|_{L^{2}(X)}\to 0.$ (30)
Since $\tilde{A}_{r}$ is compact, we have that $\tilde{A}_{r}f_{n}$ converges up to a subsequence, and as such by equation (30) so does $\left([\tilde{A}_{r}1]-r^{2}\lambda\right)f_{n}$, with limit $g\in L^{2}(X)$. Consider $\delta>0$ such that $0\leq\lambda+\delta/r^{2}<\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}}$ and define
$b_{r}(x):=[\tilde{A}_{r}1]-r^{2}\lambda\geq\delta.$ (31)
Thus we have that $f:=g/b_{r}\in L^{2}(X,\mu)$, and so
$\delta\|f_{n}-\frac{g}{b_{r}}\|_{L^{2}(X)}\leq\|b_{r}f_{n}-g\|_{L^{2}(X)}\rightarrow 0.$ (32)
Thus $f_{n}$ converges in $L^{2}(X,\mu)$ to the limit function $f$. Using continuity of $\tilde{\Delta}_{r}$ we conclude that
$\tilde{\Delta}_{r}f=\lambda f,$ (33)
and thus $\lambda$ is in the point spectrum.
To prove that $\lambda$ is an isolated point, we suppose by contradiction that there exists an infinite sequence $(\lambda_{n})\subset\sigma(-\tilde{\Delta}_{r})$ of distinct values such that $\lambda_{n}\rightarrow\lambda$. Then there exists $\delta>0$ such that for any high enough $n$,
$\lambda_{n}+\delta/r^{2}<\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}},\quad\quad\lambda_{n}\rightarrow\lambda.$ (34)
From the previous paragraph, we know that $\lambda_{n}$ is in the point spectrum, thus there exists $f_{n}\in L^{2}(X,\mu)$ satisfying $\|f_{n}\|_{L^{2}(X)}=1$, such that
$\tilde{\Delta}_{r}f_{n}=\lambda_{n}f_{n}.$ (35)
This can be written as
$\frac{\tilde{A}_{r}}{r^{2}}f_{n}=(\lambda_{n}-\frac{[\tilde{A}_{r}1]}{r^{2}})f_{n}.$ (36)
Using compactness of $\tilde{A}_{r}$, we know that $\tilde{A}_{r}f_{n}$ converges up to a subsequence. This implies that $(r^{2}\lambda_{n}-[\tilde{A}_{r}1])f_{n}$ converges up to a subsequence to some $g\in L^{2}(X,\mu)$. Define
$b_{r,n}(x):=[\tilde{A}_{r}1]-r^{2}\lambda_{n}.$ (37)
With this we have that
$\delta\left\|f_{n}-\frac{g}{b_{r}}\right\|_{L^{2}(X)}$ $\leq\delta\left(\left\|f_{n}-\frac{g}{b_{r,n}}\right\|_{L^{2}(X)}+\left\|\frac{g}{b_{r,n}}-\frac{g}{b_{r}}\right\|_{L^{2}(X)}\right)$ (38)
$\leq\|b_{r,n}f_{n}-g\|_{L^{2}(X)}+\delta\left\|\frac{g}{b_{r,n}}-\frac{g}{b_{r}}\right\|_{L^{2}(X)}.$ (39)
We have that $\|b_{r,n}f_{n}-g\|_{L^{2}(X)}\rightarrow 0$ and also $\|\frac{g}{b_{r,n}}-\frac{g}{b_{r}}\|_{L^{2}(X)}\rightarrow 0$ since $0<\delta\leq b_{r,n},b_{r}$ and $\lambda_{n}\rightarrow\lambda$. Thus $f_{n}$ converges. However since all the eigenvalues are different, we know that $\langle f_{n},f_{j}\rangle=\delta_{n,j}$, and so the sequence cannot converge up to a subsequence, achieving contradiction. This shows that $\lambda$ is an isolated point of $\sigma(-\tilde{\Delta}_{r})$ finishing the first part of the proof.
Let us now prove that $\tilde{\Delta}_{r}-\lambda I$ is a Fredholm operator.
To show that $\ker(-\tilde{\Delta}_{r}-\lambda I)$ is finite dimensional we proceed by contradiction. Assume that there exists an infinite sequence $(f_{n})\subset\ker(-\tilde{\Delta}_{r}-\lambda I)$ such that $\langle f_{n},f_{j}\rangle=\delta_{n,j}$. Thus
$\frac{\tilde{A}_{r}f_{n}}{r^{2}}=\left(\lambda-\frac{[\tilde{A}_{r}1]}{r^{2}}\right)f_{n}.$ (40)
Similar to before we can use compactness of $\tilde{A}_{r}$ and the condition on $\lambda$ to conclude that $f_{n}$ converges in $L^{2}(X,\mu)$ up to a subsequence. However, this is prevented by $\langle f_{n},f_{j}\rangle=\delta_{n,j}$.
To show that the image of $\tilde{\Delta}_{r}-\lambda I$ is closed, consider a sequence $g_{n}:=(-\tilde{\Delta}_{r}-\lambda I)(f_{n})$ such that $g_{n}\rightarrow g$. Similarly to before, we can conclude that since $g_{n}$ converges, then $f_{n}$ converges to some $f$, and thus $g=(-\tilde{\Delta}_{r}-\lambda I)(f)$. ∎
Corollary 3.13.
Let $(X,\mathsf{d},\mu)$ be a metric measure space such that for some $r_{0}>0$ the assumption ($\mathrm{II}_{r}$) holds for any $r\in(0,r_{0})$. Then
$\lim\limits_{r\downarrow 0}\big{(}\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})\big{)}=+\infty.$
Proof.
Proposition 3.12 implies that
$\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}}\leq\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r}).$ (41)
But for any $r>0$,
$\tilde{A}_{r}1=\frac{1}{2}\left(A_{r}1+A^{*}_{r}1\right)\geq\frac{1}{2}A_{r}1=\frac{1}{2}$
hence (41) implies that
$\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})\geq\frac{1}{2r^{2}}\stackrel{{\scriptstyle r\downarrow 0}}{{\to}}+\infty.$
Let us provide our second spectral result on $\tilde{\Delta}_{r}$.
Proposition 3.14.
Let $(X,\mathsf{d},\mu)$ be a connected metric measure space satisfying ($\mathrm{II}_{r}$) for some fixed $r>0$. Then the kernel of $\tilde{\Delta}_{r}$ contains constant functions only, and $\tilde{E}_{r}$ defines a scalar product on
$\mathrm{\Pi}(X,\mu)\coloneqq\left\{f\in L^{2}(X,\mu):\int_{X}f\mathop{}\!\mathrm{d}\mu=0\right\}.$ (42)
Proof.
Consider $f\in L^{2}(X,\mu)\backslash\{0\}$ such that $\tilde{\Delta}_{r}f=0$. Then $\tilde{E}_{r}(f)=0$. This implies that for $\mu$-a.e. $x\in X$,
$\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\left(\frac{f(x)-f(y)}{r}\right)^{2}d\mu(y)=0$ (43)
which implies, in turn,
$\mu\left(\{y\in B_{r}(x):f(y)=f(x)\}\right)=\mu(B_{r}(x)).$ (44)
Consider $F\coloneqq\{x\in X:\text{$f(x)$ is a well-defined real number}\}$ and
$A:=\{x\in F:\mu\left(\{y\in B_{r}(x):f(y)=f(x)\}\right)=\mu(B_{r}(x))\}.$ (45)
Then
[…]
Take $z\in A$ and let $c=f(z)$. Consider
$I=\{x\in A:f(x)=c\},\quad\quad I^{\prime}=\{x\in A:f(x)\neq c\},$ (47)
and notice that $I\cup I^{\prime}=A$. Suppose by contradiction that $I^{\prime}\neq\emptyset$. Set
$W:=\bigcup_{x\in I}B_{r}(x),\quad\quad V:=\bigcup_{y\in I^{\prime}}B_{r}(y).$ (48)
Since $V$ and $W$ are open sets whose union contains $A$ which has full measure in $X$, we must have
$W\cup V=X,$ (49)
[…]
$B_{r}(x)\cap B_{r}(y)\neq\emptyset.$ (50)
However
$f|_{B_{r}(x)\cap B_{r}(y)}(w)$ $=f(x)\quad\quad\mu\text{-a.e. }w\in X,$ (51)
$f|_{B_{r}(x)\cap B_{r}(y)}(w)$ $=f(y)\quad\quad\mu\text{-a.e. }w\in X.$ (52)
This is not possible since $f(x)\neq f(y)$ and $\mu(B_{r}(x)\cap B_{r}(y))>0$. This implies that $I=A$ and that $f(w)=c$ for $\mu$-a.e. $w\in X$. Then $\tilde{\Delta}_{r}$ has a non-trivial kernel consisting of the constant functions only. Moreover, since $\tilde{\Delta}_{r}$ is non-negative (Remark 3.8), we get the desired property on (42). ∎
We are now in a position to prove Theorem 1. We recall that the context of this statement is a compact uniformly locally doubling metric measure space $(X,\mathsf{d},\mu)$. The compactness of the space ensures that ($\mathrm{II}_{r}$) holds for any $r>0$, see Lemma 2.6.
Proof.
For $k\geq 1$ integer, let $x_{0},\ldots,x_{k}\in X$ be distinct points. Set
$\bar{r}_{k}\coloneqq\min_{0\leq i\neq j\leq k}\frac{\mathsf{d}(x_{i},x_{j})}{4}\,\cdot$
For any $i\in\{0,\ldots,k\}$ and $y\in X$, define
$\tilde{f}_{i}(y)\coloneqq\left(1-\frac{\mathsf{d}(x_{i},y)}{\bar{r}_{k}}\right)^{+}\qquad\text{and}\qquad f_{i}(y)\coloneqq\frac{\tilde{f}_{i}(y)}{\|\tilde{f}_{i}\|_{2}}\,\cdot$
Note that each $f_{i}$ is an $L^{2}$-normalized Lipschitz function supported in $B_{\bar{r}_{k}}(x_{i})$, and that $(f_{0},\ldots,f_{k})$ is an orthonormal family of $L^{2}(X,\mu)$. Set
$V\coloneqq\mathrm{Span}(f_{0},\ldots,f_{k})\in\mathcal{G}_{k+1}(L^{2}(X,\mu))$
and observe that for any $r>0$,
$\tilde{\lambda}_{k,r}\leq\max_{\begin{subarray}{c}f\in V\\ \|f\|_{2}=1\end{subarray}}\tilde{E}_{r}(f).$
Consider $r\in(0,\bar{r}_{k})$ and $i\neq j$ in $\{0,\ldots,k\}$. If $x\in B_{2\bar{r}_{k}}(x_{i})$, then $f_{j}(y)=0$ for any $y\in B_{r}(x)$, while if $x\notin B_{2\bar{r}_{k}}(x_{i})$, then $f_{i}(y)=0$ for any $y\in B_{r}(x)$. In both cases,
$(f_{i}(x)-f_{i}(y))(f_{j}(x)-f_{j}(y))=0$
for any $y\in B_{r}(x)$. Thus
$\tilde{E}_{r}(f_{i},f_{j})=0.$ (53)
Consider $f\in V$ such that $\|f\|_{2}=1$. Then $f=\sum_{i=0}^{k}a_{i}f_{i}$ for some $a_{0},\ldots,a_{k}\in\mathbb{R}$ such that $\sum_{i=0}^{k}a_{i}^{2}=1$. By (53), we get
$\tilde{E}_{r}(f)=\sum_{i=0}^{k}a_{i}^{2}\tilde{E}_{r}(f_{i})\leq\max_{0\leq i\leq k}\tilde{E}_{r}(f_{i}).$
Let $r_{0},C$ be the parameters of the uniform local doubling property of $(X,\mathsf{d},\mu)$, see Definition 2.8. Then for any $i\in\{0,\ldots,k\}$ and $r\in(0,r_{0})$,
$\tilde{E}_{r}(f_{i})$ $\leq\frac{\operatorname{Lip}^{2}(f_{i})}{4}\int_{X}\fint_{B_{r}(x)}\underbrace{\left(1+\frac{V(x,r)}{V(y,r)}\right)}_{\leq 1+C^{2}}\underbrace{\frac{\mathsf{d}^{2}(x,y)}{r^{2}}}_{\leq 1}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$\leq\frac{\operatorname{Lip}^{{2}}(f_{i})(1+C^{2})\mu(X)}{4}\,\cdot$
Therefore, for any $r<\tilde{r}_{k}\coloneqq\min(\bar{r}_{k},R)$, we get
$\tilde{\lambda}_{k,r}\leq\tilde{C}\coloneqq\frac{(1+C^{2})\mu(X)}{4}\max_{0\leq i\leq k}\operatorname{Lip}^{{2}}(f_{i})\,\cdot$ (54)
By Corollary 3.13, there exists $r_{k}\in(0,\tilde{r}_{k})$ such that $\tilde{C}<\min\sigma_{\mathrm{ess}}(-\tilde{\Delta}_{r})$ for any $r\in(0,r_{k})$. Then Remark 3.11 implies that for such an $r$ the operator $\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues $\lambda_{0}(-\tilde{\Delta}_{r})\leq\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$ such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$. That $\lambda_{0}(-\tilde{\Delta}_{r})=0$ follows from Proposition 3.14. Moreover, by Remark 3.11 and Proposition 3.14, we know that
$\tilde{\lambda}_{1,r}=\min_{f\in\Pi(X,\mu)}\frac{\tilde{E}_{r}(f)}{\|f\|_{2}}$
where $\Pi(X,\mu)$ is as in (42). Since $\Delta_{r}$ has a kernel which is $L^{2}$-orthogonal to $\Pi(X,\mu)$ (Proposition 3.14), we have $\tilde{E}_{r}(f)>0$ for any $f\in\Pi(X,\mu)$, hence we get
$\tilde{\lambda}_{1,r}>0.$
4. First eigenvalue of torus and hypercubes
In this section, we derive some results which will be applied in Section 6. We let $m$ be a positive integer kept fixed throughout the section.
4.1. A preliminary lemma
We begin with a result where we use the normalized sinc function, namely
[…]
and the following notation: for any $p=(p_{1},...,p_{m})\in\mathbb{Z}^{m}$,
$J(p)\coloneqq\{i\in\{1,\ldots,m\}:p_{i}\neq 0\},\qquad j(p)\coloneqq\#J(p).$ (55)
Lemma 4.1.
$\liminf_{r\rightarrow 0}\inf_{0\neq p\in\mathbb{Z}^{m}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|>0.$ (56)
Proof.
[…]
$\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\neq 1,$
so that
$\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|>0.$ (57)
Moreover, for any $r>0$, if $|p|_{\infty}\coloneqq\max_{1\leq i\leq m}|p_{i}|\to+\infty$, then
$\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|\rightarrow\frac{1}{r^{2}}>0,$
hence there exists $R>0$ such that
$\inf_{0\neq p\in\mathbb{Z}^{m}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|=\min_{\begin{subarray}{c}0\neq p\in\mathbb{Z}^{m}\\ |p|_{\infty}<R\end{subarray}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|>0.$
[…]
$\lim_{n}\left|\frac{1}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p^{(n)}_{i}r)\right)\right|=0.$ (58)
We have two cases.
  • There exists $\alpha>0$ and $j\in\{1,\ldots,m\}$ such that $\liminf_{n}p^{(n)}_{j}r_{n}>\alpha$.
  • For any $j\in\{1,\ldots,m\}$, one has $\liminf_{n}p^{(n)}_{j}r_{n}=0$.
If the first one were true, then we would have
$\liminf_{n}\left|1-\prod_{i\in J(p)}\operatorname{sinc}(p^{(n)}_{i}r)\right|>|1-\alpha|,$ (59)
and so (58) couldn’t hold. On the contrary, if the second case were true, up to extracting a subsequence we would have $p^{(n)}_{j}r_{n}\to 0$ for any $j\in\{1,\ldots,m\}$. For any $y=(y_{1},\ldots,y_{m})\in\mathbb{R}^{m}$, set
$G(y)\coloneqq 1-\prod_{j=1}^{m}\operatorname{sinc}(y_{j}).$ (60)
Then $G$ is smooth on $\mathbb{R}^{m}$ and satisfies
$G(y)=\frac{1}{2}|y|^{2}+o(|y|^{3}),\qquad|y|\to 0.$ (61)
As a consequence, for $y\in\mathbb{R}^{m}$ such that $|y|$ is small enough,
$\left|G(y)\right|\geq\frac{1}{4}|y|^{2}.$ (62)
Then, for large enough $n$,
$\left|1-\prod_{i\in I(p^{(n)})}\operatorname{sinc}(p^{(n)}_{i}r_{n})\right|$ $=G(r_{n}p^{(n)})\geq\frac{1}{4}|r_{n}p^{(n)}|^{2}\geq\frac{1}{4}r_{n}^{2},$
because $p^{(n)}\neq 0$ implies that there exists at least one $i$ such that $|p^{(n)}_{i}|\geq 1$. Thus we obtain
$\lim_{n}\left|\frac{1}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(\pi p^{(n)}_{i}r_{n})\right)\right|\geq\frac{1}{4}$ (63)
and so (58) could not hold. This concludes the proof. ∎
4.2. Torus
Consider the torus
$\mathbb{T}^{m}:=\mathbb{R}^{m}/\big{(}-1+2\mathbb{Z}\big{)}^{m}$
with its natural quotient map $\pi:\mathbb{R}^{m}\to\mathbb{T}^{m}$. Let $\pi^{-1}$ be the inverse of the bijective map
$\pi:[-1,1)^{m}\to\mathbb{T}^{m}.$
Let $\mathsf{d}_{\infty}$ be the distance in $\mathbb{R}^{m}$ associated with the infinity norm, given by
$\mathsf{d}_{\infty}(x,y):=\max\{|x_{i}-y_{i}|:i\in\{1,...,m\}\}$ (64)
for any $x=(x_{1},\ldots,x_{m})$ and $y=(y_{1},\ldots,y_{m})$ in $\mathbb{R}^{m}$. With respect to this distance, the open ball of radius $r>0$ centered at $x\in\mathbb{R}^{m}$ is
$Q_{r}(x):=\prod_{i=1}^{m}(-r+x_{i},x_{i}+r).$ (65)
For any $x,y\in\mathbb{T}^{m}$, set
$\tilde{\mathsf{d}}_{\infty}(x,y)=\inf_{z\in\pi^{-1}(x),w\in\pi^{-1}(y)}\mathsf{d}_{\infty}(z,w).$
Then $\tilde{\mathsf{d}}_{\infty}$ defines a distance on $\mathbb{T}^{m}$, and we denote by $\tilde{Q}_{r}(x)$ the open ball of radius $r>0$ centered at $x\in\mathbb{T}^{m}$ with respect to this distance. We also introduce the probability measure
$\mathbb{L}^{m}\coloneqq\pi_{\#}\left(\frac{\mathcal{L}^{m}}{2^{m}}\right)$
on $\mathbb{T}^{m}$. Note that this is also the normalized Haar measure of $\mathbb{T}^{m}$ seen as a Lie group.
It is obvious that the metric measure space $\mathfrak{T}^{m}\coloneqq(\mathbb{T}^{m},\tilde{\mathsf{d}}_{\infty},\mathbb{L}^{m})$ satisfies the assumptions of Theorem 1, hence we know that for any small enough $r$,
$\tilde{\lambda}_{1,r}=\lambda_{1}(-\tilde{\Delta}_{r})>0.$
Then the following holds.
Proposition 4.2.
$\liminf_{r\rightarrow 0}\lambda_{1}(-\tilde{\Delta}_{r})>0.$ (66)
Proof.
For any $x\in\mathbb{T}^{m}$ and $r>0$ small enough, the ball $\tilde{Q}_{r}(x)\subset\mathbb{T}^{m}$ is given by
$)\tilde{Q}_{r}(x)=\pi(Q_{r}(\tilde{x}$)) (67)
[…]
$\mathbb{L}^{m}(\tilde{Q}_{r}(x))=\frac{\mathcal{L}^{m}(Q_{r}(\tilde{x}))}{2^{m}}=r^{m}$ (68)
so that the $r$-energy functional of $\mathfrak{T}^{m}$ writes as
$\tilde{E}_{r,\mathfrak{T}^{m}}(f)=\frac{1}{r^{m}}\int_{\mathbb{T}^{m}}\int_{\tilde{Q}_{r}(x)}\left(\frac{f(x)-f(y)}{r}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$ (69)
for any $f\in L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$.
We act by contradiction. Assume that there exist $r_{n}\rightarrow 0$ and $\{f_{n}\}\subset L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$ satisfying $\int_{\mathbb{T}^{m}}f_{n}\mathop{}\!\mathrm{d}\mathbb{L}^{m}=0$ and $\|f_{n}\|_{L^{2}(\mathbb{T}^{m})}=1$, such that
$\tilde{\Delta}_{r_{n}}f_{n}=\lambda_{1}(-\tilde{\Delta}_{r_{n}})f_{n},\quad\quad\lambda_{1}(-\tilde{\Delta}_{r_{n}})\rightarrow 0.$ (70)
With no loss of generality, we assume that each $r_{n}$ is small enough to ensure that $\tilde{E}_{r_{n},\mathfrak{T}^{m}}$ writes as in (69).
From (68) we can write, for any $n$ and $\mathbb{L}^{m}$-a.e. $x\in\mathbb{T}^{m}$,
$-\tilde{\Delta}_{r_{n}}f_{n}(x)$ $=\frac{1}{r_{n}^{2}}\left(f_{n}(x)-\frac{1}{r_{n}^{m}}\int_{\mathbb{T}^{m}}1_{\tilde{Q}_{r_{n}}(x)}(y)f_{n}(y)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)\right)$
$=\frac{1}{r_{n}^{2}}\left(f_{n}(x)-\frac{1}{r_{n}^{m}}\bigg{(}1_{\tilde{Q}_{r_{n}}(0)}*f_{n}\bigg{)}(x)\right).$ (71)
We consider the Fourier decomposition of $f_{n}$, $1_{\tilde{Q}_{r_{n}}(0)}$, and $\tilde{\Delta}_{r_{n}}f_{n}$, namely
$f_{n}=\sum_{p\in\mathbb{Z}^{m}}a_{p,n}e_{p}$
$1_{\tilde{Q}_{r_{n}}(0)}=\sum_{p\in\mathbb{Z}^{m}}b_{p,n}e_{p}$
$\tilde{\Delta}_{r_{n}}f_{n}=\sum_{p\in\mathbb{Z}^{m}}c_{p,n}e_{p}$
where $\{e_{p}\}_{p\in\mathbb{Z}^{m}}$ is the orthonormal basis of $L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$ given by
$e_{p}:\mathbb{T}^{m}\ni x\mapsto e^{i\pi p\cdot\pi^{-1}(x)}\qquad\forall\,p\in\mathbb{Z}^{m}.$
Since the Fourier coefficients of a convolution are the product of the coefficients, we obtain from (71) that
$c_{p,n}=\frac{1}{r_{n}^{2}}\left(a_{p,n}-\frac{b_{p,n}}{r_{n}^{m}}a_{p,n}\right)=\frac{a_{p,n}}{r_{n}^{2}}\left(1-\frac{b_{k,n}}{r_{n}^{m}}\right).$ (72)
We can compute each coefficient $b_{p,n}$ by means of Fubini’s Theorem; we obtain
$b_{p,n}=\int_{Q_{r_{n}}(0)}e^{i\pi p\cdot x}\frac{d\mathcal{L}^{m}(x)}{2^{m}}=r_{n}^{m}\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n}).$ (73)
Thus
$c_{p,n}=\frac{a_{p,n}}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n})\right).$ (74)
[…]
$\frac{1}{r_{n}^{2}}\left|1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n})\right|\geq\alpha.$ (75)
This implies that
$|c_{p,n}|^{2}\geq|a_{p,n}|^{2}\alpha^{2}.$ (76)
By Parseval’s identity, and since $f_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$ we have $a_{0,n}=0$,
[…]
in contradiction with $\|-\tilde{\Delta}_{r_{n}}f_{n}\|_{L^{2}(\mathbb{T}^{m})}\rightarrow 0$ provided by (70). ∎
4.3. Shrinking hypercubes
For any $b>0$, consider the metric measure space $\mathfrak{Q}^{m}(b)\coloneqq([0,b]^{m},\mathsf{d}_{\infty},\mathcal{L}^{m})$. It trivially satisfies the assumptions of Theorem 1, so that for any small enough $r$,
$\tilde{\lambda}_{1,r}=\lambda_{1}(-\tilde{\Delta}_{r,\mathfrak{Q}^{m}(b)})>0.$ (78)
Then the following holds.
Lemma 4.3.
$\liminf_{b\rightarrow 0}\lim_{r\to 0}\lambda_{1}(-\tilde{\Delta}_{r,\mathfrak{Q}^{m}(b)})=+\infty.$ (79)
Proof.
We suppose by contradiction that (79) fails. Then there exist $b_{n}\rightarrow 0$ and $r_{n}\rightarrow 0$ such that
$\lim_{n}\lambda_{1}(-\tilde{\Delta}_{r_{n},\mathfrak{Q}^{m}(b)})<+\infty.$
Since we are first taking the limit in $r$ and then in $b$, we can assume that $\overline{r}_{n}:=\frac{r_{n}}{b_{n}}\rightarrow 0$. For any $n$, by a simple scaling argument we have
$\lambda_{1}(-\tilde{\Delta}_{r_{n},\mathfrak{Q}^{m}(b_{n})})=\frac{1}{b_{n}^{2}}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})$
thus
$\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\rightarrow 0.$ (80)
From (78), assuming that each $r_{n}$ is small enough, we know that there exists $f_{n}\in\Pi([0,1]^{m},\mathcal{L}^{m})$ such that $\|f_{n}\|_{L^{2}([0,1]^{m},\mathcal{L}^{m})}=1$ and
$\tilde{E}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}(f_{n})=\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}).$
Consider the continuous function
$\begin{array}[]{cccccc}T&:&\mathbb{T}^{m}&\rightarrow&[0,1]^{m}\\ &&x&\mapsto&(|\bar{x}_{1}|,...,|\bar{x}_{m}|)\end{array}$
where $\bar{x}=\pi^{-1}(x)\in[-1,1)^{m}$. For any $n$, set
$\tilde{f}_{n}=f_{n}\circ T\in L^{2}(\mathbb{T}^{m},\mathbb{L}^{m}).$
From this we have that $f_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$ Let us prove that
$\|\tilde{f}_{n}\|_{L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})}=1.$
Let $C_{1},\ldots,C_{2^{m}}$ denote the $2^{m}$ sets of the form $I_{1}\times\ldots\times I_{m}$ where each $I_{i}$ is either [1,0][-1,0] or $[0,1]$. For any $j\in\{1,\ldots,2^{m}\}$, set
$\begin{array}[]{cccccc}N_{j}&:&[0,1]^{m}&\rightarrow&C_{j}\\ &&(\xi_{1},\ldots,\xi_{m})&\mapsto&(\varepsilon_{i}\xi_{1},\ldots,\varepsilon_{m}\xi_{m})\end{array}$
where $\varepsilon_{i}$ is $1$ if $I_{i}=[0,1]$ and 1-1 otherwise. Note that $N_{j}$ is an isometry which preserves the Lebesgue measure, and that $T\circ\pi\circ N_{j}$ is equal to the identity. Then
$\|\tilde{f}_{n}\|_{L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})}^{2}=\int_{\mathbb{T}^{m}}(f_{n}\circ T)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}$ $=\frac{1}{2^{m}}\int_{[-1,1]^{m}}(f_{n}\circ T\circ\pi)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}$
$=\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{C_{j}}(f_{n}\circ T\circ\pi)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}$
$=\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{[0,1]^{m}}(f_{n}\circ T\circ\pi\circ N_{j})^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}$
$=\|f_{n}\|_{L^{2}([0,1]^{m},\mathcal{L}^{m})}^{2}=1.$ (81)
We claim that
$\tilde{E}_{\overline{r}_{n},\mathfrak{T}^{m}}(\tilde{f}_{n})\leq 2^{m}\tilde{E}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}(f_{n}).$ (82)
Since $\tilde{f}_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$, the latter provides a contradiction with Proposition 4.2, namely
$0<\lambda_{1}(-\tilde{\Delta}_{\bar{r}_{n},\mathfrak{T}^{m}})\leq 2^{m}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\to 0.$
Given $x\in\mathbb{T}^{m}$, define
$\tilde{G}_{n}(x)\coloneqq\frac{1}{\overline{r}_{n}^{m}}\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{\tilde{f}_{n}(x)-\tilde{f}_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y).$
For any $x\in[0,1]^{m}$, set
${G}_{n}(x)$ $\coloneqq\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{1}{V(x,\overline{r}_{n})}+\frac{1}{V(y,\overline{r}_{n})}\right)$
$\qquad\qquad\quad\times\left(\frac{f_{n}(x)-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$
where
$V(z,r)\coloneqq\mathcal{L}^{m}(Q_{r}(z)\cap[0,1]^{m})$ (83)
for any $z\in[0,1]^{m}$. Then
$4\tilde{E}_{\overline{r}_{n},\mathfrak{T}^{m}}(\tilde{f}_{n})$ $=\int_{\mathbb{T}^{m}}\tilde{G}_{n}(x)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$
$4\tilde{E}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}(f_{n})$ $=\int_{[0,1]^{m}}{G}_{n}(x)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x).$
Moreover, for all $x\in\mathbb{T}^{m}$,
$\frac{1}{\overline{r}_{n}^{m}}\leq\frac{2^{m}}{\mathcal{L}^{m}(Q_{\overline{r}_{n}}(T(x))\cap[0,1]^{m})}=\frac{2^{m}}{V(T(x),\overline{r}_{n})}.$ (84)
For any $x\in\mathbb{T}^{m}$, there exists some $k\in\{1,...,2^{m}\}$ and $\overline{x}\in C_{k}$ such that $\pi(\overline{x})=x$. We will now consider for each $j\in\{1,...,2^{m}\}$ the rectangle given by
$R_{n,j}(x):=T(\pi(C_{j})\cap\tilde{Q}_{\overline{r}_{n}}(x))\subset[0,1]^{m}.$ (85)
and we point out that
$R_{n,j}(x)\subset R_{n,k}(x)=[0,1]^{m}\cap Q_{\overline{r}_{n}}(T(x))$ (86)
for any $j\in\{1,...,2^{m}\}$. This follows since for each $i\in\{1,...,m\}$ we have 4 possibilities
  1. (1)
    $|\overline{x}_{i}|<\overline{r}_{n}$
  2. (2)
    $|1-\overline{x}_{i}|<\overline{r}_{n}$
  3. (3)
    $|-1-\overline{x}_{i}|<\overline{r}_{n}$
  4. (4)
    $\neg\left((1)\lor(2)\lor(3)\right)$.
If $\pi_{i}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the projection in the $i$-th coordinate, we conclude
[…]
Thus we conclude that for all $i\in\{1,...,m\}$ and $l\in\{1,...,2^{m}\}$ we have $\pi_{i}(R_{n,l}(x))\subset\pi_{i}(R_{n,k}(x))$, and since these sets are rectangles, we conclude equation (86).
From (84) we can deduce
$\tilde{G}_{n}(x)$ $=\frac{1}{\bar{r}_{n}^{m}}\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{\tilde{f}_{n}(x)-\tilde{f}_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (89)
$\leq\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)$ (90)
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (91)
$=\sum_{j=1}^{2^{m}}\int_{\tilde{Q}_{\overline{r}_{n}}(x)\cap\pi(C_{j})}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)$ (92)
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (93)
$=\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j}}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(\pi(y)),\overline{r}_{n})}\right)$ (94)
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(y)))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (95)
For each integral, change coordinates by $N_{j}$ to conclude
$\tilde{G}_{n}(x)$ $\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{N_{j}^{-1}(\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j})}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(\pi(N_{j}(y))),\overline{r}_{n})}\right)\times$ (96)
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(N_{j}(y))))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y).$ (97)
We have that
$N_{j}^{-1}(\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j})=R_{n,j}(x),$ (98)
so by equation (86) and the fact that $T\circ\pi\circ N_{j}=id$, we conclude
$\tilde{G}_{n}(x)$ $\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{R_{n,j}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)\times$ (99)
$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (100)
$\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{R_{n,k}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(y,\overline{r}_{n})}\right)\times$ (101)
$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (102)
$=\int_{[0,1]^{m}\cap Q_{\overline{r}_{n}}(T(x))}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(y,\overline{r}_{n})}\right)\times$ (103)
$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (104)
$=2^{m}G_{n}(T(x)).$ (105)
Now we integrate both sides in $\mathbb{T}^{m}$ and change variables by $\pi$ and $N_{j}$ to conclude
$4\tilde{E}_{\overline{r}_{n},\mathfrak{T}^{m}}(\tilde{f}_{n})=\int_{\mathbb{T}^{m}}\tilde{G}_{n}(x)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$ $\leq 2^{m}\int_{\mathbb{T}^{m}}G_{n}(T(x))\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$ (106)
$=\int_{[-1,1)^{m}}G_{n}(T(\pi(x)))\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (107)
$=\sum_{j=1}^{2^{m}}\int_{C_{j}}G_{n}(T(\pi(x)))\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (108)
$=\sum_{j=1}^{2^{m}}\int_{[0,1]^{m}}G_{n}(T(\pi(N_{j}(x))))\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (109)
$=\sum_{j=1}^{2^{m}}\int_{[0,1]^{m}}G_{n}(x)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (110)
$=2^{m}4\tilde{E}_{\overline{r}_{n},[0,1]^{m}}(f_{n})=2^{m}4\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}).$ (111)
With this we obtain (82). ∎
5. $L^{2}$ convergence
In this section, we prove Theorem 2. Let $M$ be a smooth, compact, connected manifold of dimension $m\geq 2$. Assume that $M$ is endowed with a smooth Riemannian metric $g$ and let $\mathsf{d}_{g}$ and $\mathrm{vol}_{g}$ be the associated Riemannian distance and volume measure on $M$.
In this smooth context, the function $V(\cdot,r)$ is obviously continuous for any $r>0$. Since $M$ is compact, this implies that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies ($\mathrm{II}_{r}$) and ($\mathrm{I}_{r}$). Then Lemma 3.3 applies and ensures that $\tilde{\Delta}_{r}$ is a bounded self-adjoint operator acting on $L^{2}(M,\mathrm{vol}_{g})$. The compactness of $M$ also ensures that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is locally Ahlfors regular: there exists a constant $C>1$ such that for any $x\in M$ and $r\in(0,\mathrm{diam}(M)]$,
$C^{-1}r^{m}\leq V(x,r)\leq Cr^{m}.$ (112)
Note that this condition trivially implies a uniform local doubling property for $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$.
5.1. Convergence in the sense of distributions
[…]
$J_{x}(\xi)=1+\underline{O}_{K}(|\xi|^{2})$
where for any $h>0$, the notation $\underline{O}_{K}(h)$ stands for a quantity independent on $x\in K$ whose absolute value divided by $h$ is bounded. Here $K$ is a compact subset of $M$. We write $\underline{O}$ instead of $\underline{O}_{M}$. Then the following holds.
Proposition 5.1.
Consider $f,\psi\in C^{2}(M)$. Then
$\lim_{r\rightarrow 0}\langle-\tilde{\Delta}_{r}f,\psi\rangle_{2}=C_{m}\int_{M}\langle\mathop{}\!\mathrm{d}f,\mathop{}\!\mathrm{d}\psi\rangle_{g}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}.$ (113)
Proof.
For any $x\in M$ and $r>0$, set
$\tilde{e}_{r}(f,\psi;x)\coloneqq\frac{1}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
so that
$\langle-\tilde{\Delta}_{r}f,\psi\rangle_{2}=\int_{M}\tilde{e}_{r}(f,\psi;y)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y).$
On one hand,
$|\tilde{e}_{r}(f,\psi;x)|$ $\leq\frac{1}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\frac{|f(x)-f(y)||\psi(x)-\psi(y)|}{r^{2}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$\leq\frac{1}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\frac{\operatorname{Lip}(f)\operatorname{Lip}(\psi)\mathsf{d}_{g}^{2}(x,y)}{r^{2}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$\leq\frac{\operatorname{Lip}(f)\operatorname{Lip}(\psi)}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y).$
By (112), we obtain
$|\tilde{e}_{r}(f,\psi;x)|\leq\frac{\operatorname{Lip}(f)\operatorname{Lip}(\psi)(1+C^{2})}{4}\,\cdot$ (114)
On the other hand, assume that $r$ is smaller than $\mathrm{i}_{M}(x)$, and consider $\tilde{f}\coloneqq f\circ\exp_{x}$ and $\tilde{\psi}\coloneqq\psi\circ\exp_{x}$ on $\mathbb{B}^{m}_{r}(0)$. The first-order Taylor expansion of $\tilde{f}$ and $\tilde{\psi}$ yields
$\tilde{f}(\xi)$ $=\tilde{f}(0)+(\mathop{}\!\mathrm{d}\tilde{f})_{0}(\xi)+\underline{O}_{\{0\}}(|\xi|^{2})$
$\tilde{\psi}(\xi)$ $=\tilde{\psi}(0)+(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}(\xi)+\underline{O}_{\{0\}}(|\xi|^{2}).$
Then
$\phantom{=}\int_{B_{r}(x)}(f(x)-f(y))(\psi(x)-\psi(y))\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$=\int_{\mathbb{B}^{m}_{r}(0)}(\tilde{f}(0)-\tilde{f}(\xi))(\tilde{\psi}(0)-\tilde{\psi}(\xi))J(\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$=\int_{\mathbb{B}^{m}_{r}(0)}((\mathop{}\!\mathrm{d}\tilde{f})_{0}(\xi)+\underline{O}_{\{0\}}(r^{2}))((\mathop{}\!\mathrm{d}\tilde{\psi})_{0}(\xi)+\underline{O}_{\{0\}}(r^{2}))(1+\underline{O}_{\{0\}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$=\sum_{i,j=1}^{m}[(\mathop{}\!\mathrm{d}\tilde{f})_{0}]_{i}[(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}]_{j}\int_{\mathbb{B}^{m}_{r}(0)}\xi_{j}\xi_{i}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{\{0\}}(r^{m+3})$
$=(\mathop{}\!\mathrm{d}\tilde{f})_{0}\cdot(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}\int_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{\{0\}}(r^{m+3}).$
Moreover, it is known that (see e.g. [MT23, Remark 2.11])
$1+\frac{V(x,r)}{V(y,r)}={2}+\underline{O}_{\{x\}}(r^{2})$
and since $V(x,r)/\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0)))\to 1$ as $r\downarrow 0$, we obtain
$\phantom{=}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)(f(x)-f(y))(\psi(x)-\psi(y))\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$=\frac{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0))({2}+\underline{O}_{\{x\}}(r^{2}))}{V(x,r)}\fint_{\mathbb{B}^{m}_{r}(0)}(\tilde{f}(0)-\tilde{f}(\xi))(\tilde{\psi}(0)-\tilde{\psi}(\xi))J(\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$=(2+\underline{O}_{\{x\}}(r^{2}))\left((\mathop{}\!\mathrm{d}\tilde{f})_{0}\cdot(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}\fint_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{\{0\}}(r^{3})\right).$
Since $(\mathop{}\!\mathrm{d}\tilde{f})_{0}\cdot(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}=\langle\mathop{}\!\mathrm{d}f,\mathop{}\!\mathrm{d}\psi\rangle_{g}(x)$ and
$\fint_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)=2r^{2}C_{m}$
by change of variable $\xi\leftrightarrow\eta/r^{2}$, we eventually obtain that
$\tilde{e}_{r}(f,\psi;x)=C_{m}\langle\mathop{}\!\mathrm{d}f,\mathop{}\!\mathrm{d}\psi\rangle_{g}(x)+\underline{O}_{\{x\}}(r)\qquad\text{as $r\to 0$.}$ (115)
By (114) and the compactness of $M$, we can apply the dominated convergence theorem to the functions $\tilde{e}_{r}(f,\psi;\cdot)$. Then we get (113). ∎
Using integration by parts in (113), we immediately obtain the following.
Corollary 5.2.
Let $\partial g$ be the Riemannian metric induced by $g$ on $\partial M$. For any $f\in\mathcal{C}^{\infty}(M)$, the following convergence holds in the sense of distributions as $r\downarrow 0$:
$(\tilde{\Delta}_{r}f)\,\mathrm{vol}_{g}\to C_{m}\bigg{(}(\Delta_{g}f)\,\mathrm{vol}_{g}+(\partial_{\nu}^{g}f)\,\mathrm{vol}_{\partial g}\bigg{)}.$
5.2. Pointwise convergence
We aim to prove Theorem 2 in a similar way as Proposition 5.1, that is to say, by means of the dominated convergence theorem. To this aim, we first establish that pointwise convergence holds $\mathrm{vol}_{g}$-a.e. on $M$. We recall that $\partial M$ is a $\mathrm{vol}_{g}$-negligible subset of $M$.
Proposition 5.3.
Let $f\in C^{\infty}(M)$. Then for any $x\in M-\partial M$,
$\lim_{r\rightarrow 0}\tilde{\Delta}_{r}f(x)=C_{m}\Delta_{g}f(x).$ (116)
Moreover, the convergence is uniform on any compact subset of $M-\partial M$.
Proof.
Let $K$ be a compact subset of $M-\partial M$. Consider $x\in K$ and $r\in(0,\mathrm{i}_{M}(x))$. Set $\tilde{f}_{x}\coloneqq f\circ\exp_{x}$. Acting like in the proof of Proposition 5.1, we get
$\tilde{\Delta}_{r}f(x)=\frac{(2+\underline{O}_{K}(r^{2}))}{2r^{2}}\fint_{\mathbb{B}^{m}_{r}(0)}\left(\tilde{f}_{x}(\xi)-\tilde{f}_{x}(0)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi).$ (117)
The second-order Taylor expansion of $\tilde{f}_{x}$ yields
$\tilde{f}_{x}(\xi)$ $=\tilde{f}_{x}(0)+(\mathop{}\!\mathrm{d}\tilde{f}_{x})_{0}(\xi)+\frac{1}{2}(\mathop{}\!\mathrm{d}^{(2)}\tilde{f}_{x})_{0}(\xi,\xi)+\underline{O}_{K}(|\xi|^{3})$
hence we get
$\fint_{\mathbb{B}^{m}_{r}(0)}\tilde{f}_{x}(\xi)-\tilde{f}_{x}(0)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$ $=\fint_{\mathbb{B}^{m}_{r}(0)}(\mathop{}\!\mathrm{d}\tilde{f}_{x})_{0}(\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$+\frac{1}{2}\fint_{\mathbb{B}^{m}_{r}(0)}(\mathop{}\!\mathrm{d}^{(2)}\tilde{f}_{x})_{0}(\xi,\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{K}(r^{3}).$
The first term vanishes by symmetry. The second term is equal to
$\frac{1}{2}\Delta\tilde{f}_{x}(0)\fint_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)=\Delta_{g}f(x)r^{2}C_{m}.$
In the end we get
$\tilde{\Delta}_{r}f(x)$ $=\frac{2+\underline{O}_{K}(r^{2})}{2r^{2}}\left(\Delta_{g}f(x)r^{2}C_{m}+\underline{O}_{K}(r^{3})\right)$
$=(1+\underline{O}_{K}(r^{2}))\left(C_{m}\Delta_{g}f(x)+\underline{O}_{K}(r)\right)$
from which follows the desired result, by letting $r\downarrow 0$. ∎
5.3. Uniform bound
We wish now to provide a uniform $L^{\infty}$ bound for the functions $\tilde{\Delta}_{r}\psi$ where $r$ is in a neighborhood of zero.
Let us first consider the case $\partial M=\emptyset$. From Proposition 5.3, we have uniform convergence
$\|\tilde{\Delta}_{r}f-C_{m}\Delta_{g}f\|_{\infty}\to 0$
so that
$\|\tilde{\Delta}_{r}f-C_{m}\Delta_{g}f\|_{2}\leq\|\tilde{\Delta}_{r}f-C_{m}\Delta_{g}f\|_{\infty}\mathrm{vol}_{g}(M)\to 0.$
Let us now deal with the case $\partial M\neq\emptyset$.
Proposition 5.4.
Assume that $\partial M\neq\emptyset$. Consider $f\in C^{\infty}_{\nu}(M)$. Then there exists $r_{0}>0$ such that
$\sup_{0<r<r_{0}}\|\tilde{\Delta}_{r}f\|_{L^{\infty}(M)}<+\infty.$ (118)
Proof.
Since $\partial M$ is compact, we can find a finite collection of smooth parameterizations $\psi_{i}:(-4,4)^{m-1}\rightarrow\partial M$ such that
$\partial M=\bigcup_{i}\psi_{i}\left(\left[-1,1\right]^{m-1}\right).$ (119)
Step 1.
We work with any of the previous $\psi_{i}$ which we denote by $\psi$. For any $x\in\partial M$, let $\nu_{x}\in T_{x}M$ be the unit inner normal vector of $\partial M$ at $x$. Since $\partial M$ is smooth, there exists $\epsilon>0$ such that the map $E:\partial M\times[0,\epsilon]\rightarrow M$ given by
$E(x,t)=\exp_{x}^{M}(t\nu_{x})$ (120)
is an embedding, and there exists a smooth family of metrics $\{g_{t}\}_{t\in[0,\varepsilon]}$ on $\partial M$ such that for any $(x,t)\in\partial M\times[0,\epsilon]$,
$(E^{*}g)_{(x,t)}=(g_{t}\oplus d\tau^{2})_{(x,t)}.$ (121)
Pulling back each metric $g_{t}$ by $\psi$ we have, for any $\xi\in(-4,4)^{m-1}$ and $v,w\in\mathbb{R}^{m-1}$,
$(\psi^{*}g_{t})_{\xi}(v,w)=v^{T}\cdot A_{(\xi,t)}\cdot w,$ (122)
for some positive definite, symmetric (m1)(m-1)-square matrix $A_{(\xi,t)}$. From the non-degeneracy of the metric and a Lipschitz bound, we have that there exist $C,\tilde{c}>0$ such that for all $t,s\in[0,\epsilon],\xi\in[-3,3]^{m-1},v\in\mathbb{R}^{m}$ we have
$\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)\geq\tilde{c}|v|^{2},$ (123)
$\left|\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)\right|\leq C|v|^{2}|t-s|.$ (124)
Claim 1.
There exists $K>0$ such that for any $t,s\in[0,\epsilon]$, $\xi\in[-3,3]^{m-1},$ $v\in\mathbb{R}^{m}$ such that
$\left|O(\xi,t,s,v)\right|\leq K|v|\cdot|t-s|,$ (125)
where
$O(\xi,t,s,v)=\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v).$ (126)
Proof.
[…]
$\left|O(\xi,t,s,v)\right|$ $=\left|\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)\right|$ (127)
$\leq M\left|\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)\right|$ (128)
$\leq MC|v|^{2}|t-s|=MC|t-s|.$ (129)
Step 2.
Let $T(\epsilon)\coloneqq E(\partial M\times[0,\epsilon])$ be the $\epsilon$ tubular neighborhood of the boundary. Then $E$ is a diffeomorphism between $\partial M\times[0,\epsilon]$ and $T(\epsilon)$. For fixed $s\in[0,\epsilon]$, consider the product metric $g_{s}\oplus d\tau^{2}$ in $\partial M\times[0,\epsilon]$ and define the metric in $T(\epsilon)$
$\eta_{s}\coloneqq(E^{-1})^{*}(g_{s}\oplus d\tau^{2}).$ (130)
Let $\mathsf{d}_{s}$ be the distance induced from this metric. Consider the map $\Phi:(-4,4)^{m-1}\times[0,\epsilon]\rightarrow M$ given by
$\Phi(\xi,t)=E(\psi(\xi),t)$ (131)
and note that (121) implies that for any $(\xi,t)\in(-4,4)^{m-1}\times[0,\epsilon]$,
$(\Phi^{*}g)_{(\xi,t)}=(\psi^{*}g_{t}\oplus\mathop{}\!\mathrm{d}\tau^{2})_{(\xi,t)}.$ (132)
[…]
Claim 2.
Let $\tilde{c}>0$ be given by (123). For any $r>0$ and $s\in[0,\varepsilon]$, if a couple $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon]$ is such that $\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-3,3]^{m-1}\times[0,\epsilon]$, then the following holds for any $y\in M$.
  1. (1)
    If $\mathsf{d}(\Phi(\xi,t),y)<r$, then the image of any $\mathsf{d}$-minimizing geodesic $\gamma:[0,1]\rightarrow M$ is contained in $\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$ and $\tilde{\gamma}=\Phi^{-1}\circ\gamma$ satisfies $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$.
  2. (2)
    If $\mathsf{d}_{s}(\Phi(\xi,t),y)<r$, then the image of any $\mathsf{d}_{s}$-minimizing geodesic $\gamma:[0,1]\rightarrow M$ is contained in $\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$ and $\tilde{\gamma}=\Phi^{-1}\circ\gamma$ satisfies $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$.
Proof.
We prove the first result only since the proof of the second one follows from similar lines. Consider a $\mathsf{d}$-minimizing geodesic $\gamma:[0,1]\rightarrow M$ from $\Phi(\xi,t)$ to $y$. Set
$\delta$ $\coloneqq\sup\left\{t\in[0,1]\,:\,\gamma(s)\in\Phi(\mathbb{B}^{m}_{r/\sqrt{c}}(\xi,t)\cap\mathbb{H}^{m})\text{ for any }s\in[0,t)\right\},$
$\tilde{\gamma}$ $\coloneqq\Phi^{-1}\circ\left.\gamma\right|_{[0,\delta]},$
and observe that $\gamma([0,1])\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{c}}(\xi,t)\cap\mathbb{H}^{m})$ if and only if $\delta=1$. We claim that
$L(\tilde{\gamma})\leq\frac{\mathsf{d}_{g}(\Phi(\xi,t),y)}{\sqrt{\tilde{c}}}\,\cdot$ (133)
Indeed, setting $(\tilde{\alpha},\tilde{\gamma}_{m})\coloneqq\tilde{\gamma}$ where $\tilde{\alpha}:[0,\delta]\rightarrow[-3,3]^{m-1}$ and $\tilde{\gamma}_{m}:[0,\delta]\rightarrow[0,\epsilon]$, we have
$\mathsf{d}(\Phi(\xi,t),y)$ $=\int_{0}^{1}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (134)
$\geq\int_{0}^{\delta}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (135)
$=\int_{0}^{\delta}(\Phi^{*}g)_{\tilde{\gamma}(w)}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w$ (136)
[…]
Now we claim that:
$\delta<1\qquad\Rightarrow\qquad\mathsf{d}_{g}(\Phi(\xi,t),y)\geq r.$ (140)
Indeed, if $\delta<1$, since $\tilde{\gamma}(0)=(\xi,t)$ and $\tilde{\gamma}(\delta)\in\partial\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}$, then
$L(\tilde{\gamma})\geq\frac{r}{\sqrt{\tilde{c}}}$ (141)
and we get $\mathsf{d}_{g}(\Phi(\xi,t),y)\geq r$ from (133). Therefore, if $\mathsf{d}_{g}(\Phi(\xi,t),y)<r$, then (140) implies that $\delta=1$ which means $\gamma([0,1])$ is included in $\Phi(\mathbb{B}^{m}_{r/\sqrt{c}}(\xi,t)\cap\mathbb{H}^{m})$, and (133) yields $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$ as desired. ∎
Claim 3.
There exists $K>0$ such that for all $(\xi,t),(\eta,s)\in[-2,2]^{m-1}\times[0,\epsilon]$ and $r>0$ such that $\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-3,3]^{m-1}\times[0,\epsilon]$:
$\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))<r\quad\Rightarrow\quad\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))<r+Kr^{2},$ (142)
$\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))\geq r\quad\Rightarrow\quad\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))\geq r-Kr^{2}.$ (143)
Proof.
Suppose that $\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))<r$. Let $\gamma:[0,1]\rightarrow M$ be a $\mathsf{d}$-minimizing geodesic between $\Phi(\xi,t)$ and $\Phi(\eta,s)$. Then by Claim 2, we have that $\gamma([0,1])\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$, and by defining $(\tilde{\alpha},\tilde{\gamma}_{m})=\tilde{\gamma}\coloneqq\Phi^{-1}\circ\gamma$, we have that $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$. Thus we obtain
$\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))$ $=\int_{0}^{1}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (144)
$=\int_{0}^{1}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w-\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (145)
$\quad+\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (146)
$=\int_{0}^{1}\left(\psi^{*}g_{\tilde{\gamma}_{m}(w)}\oplus d\tau^{2}\right)_{(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w))}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w$ (147)
$\quad-\int_{0}^{1}\left(\psi^{*}g_{t}\oplus d\tau^{2}\right)_{(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w))}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w$ (148)
$\quad+\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (149)
$=\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (150)
$\quad+\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w,$ (151)
where we use (126) to get the last equality. By Claim 1, we have that
$\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w\right|$ $\leq\int_{0}^{1}\left|O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\right|\mathop{}\!\mathrm{d}w$ (152)
$\leq\int_{0}^{1}K\left|\dot{\tilde{\gamma}}(w)\right|\left|t-\tilde{\gamma}_{m}(w)\right|\mathop{}\!\mathrm{d}w.$ (153)
By Claim 2, we have that $|t-\tilde{\gamma}_{m}(s)|<\frac{r}{\sqrt{\tilde{c}}}$ and $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$, hence we get
$\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w\right|$ $\leq K\frac{r}{\sqrt{\tilde{c}}}\int_{0}^{1}|\dot{\tilde{\gamma}}(w)|\mathop{}\!\mathrm{d}w<K\frac{r^{2}}{\tilde{c}}\,\cdot$ (154)
Thus
$\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))$ $\leq\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (155)
$\leq\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))+\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w\right|$ (156)
$\leq\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))+K\frac{r^{2}}{\tilde{c}}$ (157)
$<r+K\frac{r^{2}}{\tilde{c}}\,\cdot$ (158)
where we use (151) to get the second inequality and (154) to get the third one. This proves (142).
To prove (143), we may assume $\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))\geq r$ and $\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))<r$. Let $\gamma:[0,1]\rightarrow M$ be a geodesic in the metric $g_{t}\oplus d\tau^{2}$. By Claim 2, we have $\gamma([0,1])\subset\Phi(\mathbb{B}^{m}_{\frac{r}{\sqrt{\tilde{c}}}}(\xi,t)\cap\mathbb{H}^{m})$ with $(\tilde{\alpha},\tilde{\gamma}_{m})=\tilde{\gamma}=\Phi^{-1}\circ\gamma$ satisfying $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$. The same estimates as before are satisfied, hence we conclude:
$r$ $<\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))\leq\int_{0}^{1}g^{\frac{1}{2}}_{\gamma(w)}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (159)
$\leq\int_{0}^{1}(\eta_{t})^{\frac{1}{2}}_{\gamma(w)}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w+\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w\right|$ (160)
$\leq\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))+\frac{K}{\tilde{c}}r^{2}.$ (161)
We omit the proof of the next elementary claim.
Claim 4.
Let $(x,t),(y,\tau)\in\partial M\times[0,\epsilon]$ and $s\in[0,\epsilon]$. Let $\gamma:[0,1]\rightarrow\partial M$ be the geodesic between $x$ and $y$ in the metric $g_{s}$. Then the curve
$w\in[0,1]\mapsto(\gamma(w),(1-w)t+w\tau)$ (162)
is a geodesic in the metric $g_{s}\oplus d\tau^{2}$.
Step 3.
For every $t\in[0,\epsilon]$, consider the exponential map $\exp^{g_{t}}$ given by the metric $g_{t}$. Since $g_{t}$ varies smoothly with respect to $t$, and $\partial M\times[0,\epsilon]$ is compact, we know that there exists $\delta>0$ lower than the injectivity radius of each $\exp^{g_{t}}$. For any $(\xi,t,\zeta,s)$ in
$D_{\delta}:=[-1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$
define
$\Psi(\xi,t,\zeta,s)\coloneqq E(\exp^{g_{t}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta),s).$ (163)
We may write
$\Psi_{(\xi,t)}(\zeta,s)=\Psi(\xi,t,\zeta,s)$
to see $\Psi$ as a function of the two last variables only, the two first being frozen. Observe that for any $(\xi,t,\zeta)$ in $1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)$,
$(g_{t})^{\frac{1}{2}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta,(d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ $=(\psi^{*}g_{t})_{\xi}^{\frac{1}{2}}(A_{(\xi,t)}^{-\frac{1}{2}}\zeta,A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ (164)
[…]
$=|\zeta|<\delta.$ (166)
Since $\delta$ is lower than the injectivity radius of the exponentials, the map
$\zeta\in\mathbb{B}^{m-1}_{\delta}(0)\mapsto\exp^{g_{t}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ (167)
is injective. Thus for every $(\xi,t)\in[-1,1]^{m-1}\times[0,\epsilon]$ the map $\Psi_{(\xi,t)}$ defined on $\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$ is a local parametrization of $M$. Moreover,
$\Psi_{(\xi,t)}(0,t)=E(\psi(\xi),t)=\Phi(\xi,t),$ (168)
and
$\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(0,t)}\right)=1.$ (169)
Claim 5.
Consider $(\xi,t)\in[-1,1]^{m-1}\times[0,\epsilon]$, and $(\zeta,s)\in\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$. Then
$\mathsf{d}_{t}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))=\sqrt{|\zeta|^{2}+(t-s)^{2}}.$ (170)
Proof.
By Claim 4, we know that the geodesic between $E^{-1}(\Psi_{(\xi,t)}(0,t))$ and $E^{-1}(\Psi_{(\xi,t)}(\zeta,s))$ in the metric $g_{s}\oplus d\tau^{2}$ is
$\gamma:w\in[0,1]$ $\mapsto(\exp^{g_{t}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}w\zeta),(1-w)t+ws)$
$\eqqcolon(\tilde{\gamma}(w),\gamma_{m}(w)).$
Then we also know that, for any $w\in[0,1]$,
$(g_{t})_{\tilde{\gamma}(w)}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))=|\zeta|.$ (171)
Thus
$\mathsf{d}_{t}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))$ $=\int_{0}^{1}\sqrt{(g_{t})_{\tilde{\gamma}(w)}^{2}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))+(s-t)^{2}}\mathop{}\!\mathrm{d}w$ (172)
$=\sqrt{|\zeta|^{2}+(s-t)^{2}}.$ (173)
Claim 6.
There exist $r_{0},\kappa>0$ such that for all $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$ and $r\in(0,r_{0})$ such that $\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\subset[-3,3]^{m-1}\times[0,\epsilon]$, we have
$\mathcal{L}^{m}\left(\mathbb{B}^{m}_{r}(0,t)\,\triangle\,\bigg{(}\Psi_{(\xi,t)}^{-1}\left(B_{r}(\Psi_{(\xi,t)}(0,t))\right)\right)\bigg{)}\leq\kappa r^{m+1}.$ (174)
Proof.
For any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$, there exists $r_{0}(\xi,t)>0$ small enough such that
$B_{r_{0}}(\Psi_{(\xi,t)}(0,t))=B_{r_{0}}(\Phi(\xi,t))\subset\Psi_{(\xi,t)}(\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]).$ (175)
By compactness of $2,2]^{m-1}\times[0,\epsilon/2]$ and continuity of the maps $\Psi_{(\xi,t)}^{-1}$, we get that there exists a common $r_{0}>0$ such that the previous holds for any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$. Consider $r<r_{0}$ and $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$, then
$B_{r}(\Psi_{(\xi,t)}(0,t))\subset\text{Im}(\Psi_{(\xi,t)}).$ (176)
Set $A_{1}:=\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))$ and $A_{2}:=\mathbb{B}^{m}_{r}(0,t).$ We will show that there exists $K>0$ such that
$A_{1}\backslash A_{2}\subset\mathbb{B}^{m}_{r+Kr^{2}}(0,t)\backslash\mathbb{B}^{m}_{r}(0,t).$ (177)
For $(\zeta,s)\in A_{1}\backslash A_{2}$, we know that
$\mathsf{d}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))<r.$ (178)
Therefore, from (168) and Claim 3, we conclude that there exists $K>0$ such that
$\mathsf{d}_{t}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))<r+Kr^{2}.$ (179)
Then we get from Claim 5 that
$\sqrt{|\zeta|^{2}+(t-s)^{2}}<r+Kr^{2}$ (180)
hence (177) is proved. A similar proof shows that
$A_{2}\backslash A_{1}\subset\mathbb{B}^{m}_{r}(0,t)\backslash\mathbb{B}^{m}_{r-Kr^{2}}(0,t).$
From the latter and (177) we conclude that
$A_{1}\triangle A_{2}\subset\mathbb{B}^{m}_{r+Kr^{2}}(0,t)\backslash\mathbb{B}^{m}_{r-Kr^{2}}(0,t).$ (181)
Thus
$\mathcal{L}^{m}(A_{1}\triangle A_{2})$ $\leq\mathcal{L}^{m}(\mathbb{B}^{m}_{r+Kr^{2}}(0,t)\backslash\mathbb{B}^{m}_{r-Kr^{2}}(0,t))$ (182)
$\leq r^{m}\mathcal{L}^{m}(\mathbb{B}^{m}_{1+Kr}(0)\backslash\mathbb{B}^{m}_{1-Kr}(0))$ (183)
$=r^{m}\omega_{m}((1+Kr)^{m}-(1-Kr)^{m})$ (184)
$\leq r^{m}C(Kr)\qquad\qquad\qquad\qquad\qquad\qquad\text{for some $C>0$}$ (185)
$\leq(CK)r^{m+1}.$ (186)
Claim 7.
There exist $C,r_{0}>0$ such that for all $r\in(0,r_{0})$ and
$(\xi,t,\zeta,s)\in[-1,1]^{m-1}\times[0,\epsilon/4]\times\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$
such that $\mathbb{B}^{m}_{2r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-2,2]^{m-1}\times[0,\epsilon/2]$ and $\mathsf{d}(\Psi_{(\xi,t)}(\zeta,s),\Phi(\xi,t))<r$, then
$\left|\frac{1}{V(\Psi_{(\xi,t)}(\zeta,s),r)}-\frac{1}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\right|\leq\frac{C}{r^{m-1}}\,\cdot$ (187)
Proof.
Let us first consider the map $G:D_{\delta}\rightarrow\mathbb{R}$ given by
$G(\xi,t,\zeta,s)\coloneqq\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(\zeta,s)}\right)^{\frac{1}{2}}.$ (188)
This map is $C^{\infty}$, and its value at any $(\xi,t)\times(0,t)$ is $1$, thus by a Taylor expansion in the variable $(\zeta,s)$ centered at $(0,t)$, and compactness of $D_{\delta}$, we obtain that there exists $k>0$ such that for all $(\xi,t,\zeta,s)\in D_{\delta}$ we have
$\left|\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(\zeta,s)}\right)^{\frac{1}{2}}-1\right|\leq k|(\zeta,s)-(0,t)|.$ (189)
In particular there exists $C>0$ such that for all $(\xi,t,\zeta,s)\in D_{\delta}$,
$\left|\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(\zeta,s)}\right)^{\frac{1}{2}}\right|\leq C.$ (190)
Let us now consider $r,\xi,t,\zeta,s$ as in the statement of the claim. Set $y\coloneqq\Psi_{(\xi,t)}(\zeta,s)$. Since $\mathsf{d}(y,\Phi(\xi,t))<r$, and $\mathbb{B}^{m}_{2r/\sqrt{\tilde{c}}}(\xi,t)\subset[-2,2]^{m-1}\times[0,\epsilon/2]$, we know by Claim 2 that $y\in\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$. Thus if we set
$(\eta,\tau)\coloneqq\Phi^{-1}(y)\in[-2,2]^{m-1}\times[0,\epsilon/2],$ (191)
we obtain that
$\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\eta,\tau)\cap\mathbb{H}^{m}\subset\mathbb{B}^{m}_{2r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-2,2]^{m-1}\times[0,\epsilon/2].$ (192)
Thus by Claim 2, we conclude that $B_{r}(y)=B_{r}(\Phi(\eta,\tau))\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\eta,\tau))$.
By (168) and (131), we easily see that
$s=\tau.$ (193)
Moreover, by (191) and (193), we also have that
$\Psi_{(\eta,s)}(0,s)=E(\psi(\eta),s)=\Phi(\eta,s)=y.$ (194)
Choose $r_{0}$ such that $\frac{r_{0}}{\sqrt{\tilde{c}}}<\frac{\epsilon}{4}$. Since $t\leq\frac{\epsilon}{4}$ and $(\eta,s)\in\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)$, this implies that $s\leq\frac{\epsilon}{2}$. Thus we can use Claim 6 to ensure that
$\mathcal{L}^{m}\left(\mathbb{B}^{m}_{r}(0,s)\triangle\left(\Psi_{(\eta,s)}\right)^{-1}\left(B_{r}(\Psi_{(\eta,s)}(0,s))\right)\right)\leq\kappa r^{m+1}.$ (195)
Then
$V(y,r)$ $=\int_{\Psi_{(\eta,z)}^{-1}(B_{r}(y))}\left|\det\left[\Psi_{(\eta,s)}^{*}g\right]_{w}\right|^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(w)$ (196)
[…]
that is, there exists $\tilde{C}>0$ such that
$\left|\mathrm{vol}_{g}(B_{r}({y}))-\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})\right|\leq\tilde{C}r^{m+1}$ (199)
Thus using the local Ahlfors regularity of $(M,g)$ and Claim 6, we obtain
$\left|\frac{1}{\mathrm{vol}_{g}(B_{r}(\overline{y}))}-\frac{1}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})}\right|$ $=\left|\frac{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})-\mathrm{vol}_{g}(B_{r}(\overline{y}))}{\mathrm{vol}_{g}(B_{r}(\overline{y}))\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})}\right|$ (200)
$\leq\frac{C\tilde{C}r^{m+1}}{r^{m}\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})}$ (201)
$\leq\frac{C\tilde{C}}{\mathcal{L}^{m}(\mathbb{B}^{m}_{1}(0)\cap\mathbb{H}^{m})r^{m-1}}$ (202)
concluding the proof. ∎
Step 4.
Claim 8.
For all $f\in C^{\infty}(M)$, there exists $C>0$ such that for all $(\xi,t,\zeta,s)\in D_{\delta}$ we have
$\left|f\circ\Psi_{(\xi,t)}(\zeta,s)-f\circ\Psi_{(\xi,t)}(0,t)\right|\leq C\|(\zeta,s)-(0,t)\|_{2},$ (203)
$\left|f\circ\Psi_{(\xi,t)}(\zeta,s)-f\circ\Psi_{(\xi,t)}(0,t)-\nabla(f\circ\Psi_{(\xi,t)})_{(0,t)}\cdot\left((\zeta,s)-(0,t)\right)\right|\leq C\|(\zeta,s)-(0,t)\|_{2}^{2}$ (204)
Also if $\partial_{\nu}f|_{\partial M}=0$, then there exists $C>0$ such that
$\left|\partial_{m}f\circ\Psi_{(\xi,t)}(0,t)\right|\leq Ct$ (205)
Proof.
The map $\tilde{f}((\xi,t),(\zeta,s))=f\circ\Psi_{(\xi,t)}(\zeta,s)$ is $C^{\infty}$, thus by a Taylor expansion of order $1$ and $2$ respectively, and compactness of $D_{\delta}$, we conclude the first two inequalities. For the last, we notice that
$\partial_{m}f\circ\Psi_{(\xi,0)}(0,0)=(\partial_{\nu}f)(\psi(\xi,0))=0.$ (206)
Thus by a Taylor expansion and compactness of $D_{\delta}$ we conclude (205). ∎
Now we will fix a function $f\in C^{\infty}(M)$ such that $\partial_{\nu}f|_{\partial M}=0$, and show that for $x\in M$ such that $d(x,\partial M)<\frac{\epsilon}{4}$ then $\Delta_{r}f(x)$ is uniformly bounded. The proof for points $x$ with $d(x,\partial M)\geq\frac{\epsilon}{4}$, follows from the uniform convergence obtained in Proposition 5.3. We will study the following term of the AMV:
$G_{r}(x):=\frac{1}{r^{2}}\int_{B_{r}(x)}\frac{1}{\mathrm{vol}_{g}(B_{r}(y))}\left(f(y)-f(x)\right)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$ (207)
since the bound for the remainder follows similarly.
We notice that by equation (119) we have that there exists some $i\in\{1,...,l\}$ and $(\xi,t)\in[-1,1]^{m-1}\times[0,\frac{\epsilon}{4}]$ such that $x=\Phi_{i}(\xi,t)$. We let $\Phi=\Phi_{i}$. Also by (175), for $0<r<r_{0}$ in the conditions of the Claim, we have that $B_{r}(x)=B_{r}(\Psi_{(\xi,t)}(0,t))\subset\Psi_{(\xi,t)}(\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon])$. Also we can choose $r_{0}$ small enough so that for all $(\xi,t)\in[-1,1]^{m-1}\times[0,{\epsilon}/{4}]$ we have $\mathbb{B}_{2r_{0}/\sqrt{\tilde{c}}}(x,t)\cap\mathbb{H}^{m}\subset[-2,2]^{m-1}\times[0,\epsilon/2]$. With this we can apply Claim 2 to conclude that $B_{r}(\Psi_{(\xi,t)(0,t)})\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t))$ and we can also apply Claim 7 for points $\Psi_{(\xi,t)}(\zeta,s)\in B_{r}(\Psi_{(\xi,t)}(0,t))$.
Thus we can change variables of the integral to obtain
$G_{r}(x)=\frac{1}{r^{2}}\int_{\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))}\frac{1}{\mathrm{vol}_{g}(B_{r}(\Psi_{(\xi,t)}(\zeta,s)))}$ (208)
$\times\left(f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))\right)\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)$ (209)
$\begin{split}G_{r}(x)&=\frac{1}{r^{2}}\int_{\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))}\frac{f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by 7 and 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by 6 and 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by LABEL:lem:determinantLemma and 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\sum_{i=1}^{m-1}\partial_{j}(f\circ\Psi_{(\xi,t)})(0,t)\zeta_{i}+\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\sum_{i=1}^{m-1}\partial_{j}(f\circ\Psi_{(\xi,t)})(0,t)\zeta_{i}+\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s))+O(1)\end{split}$ (by 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s))+O(1)\end{split}$ (by symmetry)
Now we separate further in two cases. First if $\mathsf{d}(x,\partial M)>2r$ then $t>2r$ and so for all $(\zeta,s)\in\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}$ we have $\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})=\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0))$, and so we conclude that
$\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s)$ (211)
$=\frac{1}{r^{2}\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0))}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s)$ (212)
$=0.$ (213)
This shows that $G_{r}(x)=O(1)$ for $x$ such that $2r<d(x,\partial M)<\frac{\epsilon}{4}$ since there are only a finite number of parametrizations. On the other hand if $d(x,\partial M)\leq 2r$, then we have by Lemma 8 that for $t\leq 2r$ then $\left|\partial_{m}(f\circ\Psi_{(x,t)})(0,t)\right|\leq 2Cr$, and so
$\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\partial_{m}(f\circ\Psi_{(x,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(y)\mathop{}\!\mathrm{d}\mathcal{L}(s)=O(1),$ (214)
which shows that $G_{r}(x)=O(1)$ for $x$ such that $d(x,\partial M)<2r$. ∎
6. Spectral convergence
In this section, we prove Theorem 3. We consider a smooth, compact, connected manifold $M^{m}$ endowed with a smooth Riemannian metric $g$. We let $\mathsf{d}_{g}$ and $\mathrm{vol}_{g}$ be the associated Riemannian distance and volume measure on $M$, respectively. If $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$), we let $\{\mu_{k}\}_{k\in\mathbb{N}}$ be the sequence of Laplace (resp. Neumann) eigenvalues of $(M,g)$.
6.1. Existence of limit eigenfunctions
Recalling that $C^{\infty}_{\nu}(M)$ is defined in (5), we define the Hilbert space
$H\coloneqq\overline{C^{\infty}_{\nu}(M)}^{\|\cdot\|_{W^{2,2}}}.$
We let $\Pi(M,\mathrm{vol}_{g})$ be defined as in (42) and we consider the operator $T:\Pi(M,\mathrm{vol}_{g})\rightarrow H^{*}$ which maps any $f\in\Pi(M,\mathrm{vol}_{g})$ to
$T(f)\coloneqq\left(v\in H\mapsto-\int_{M}f\Delta_{g}v\mathop{}\!\mathrm{d}\mathrm{vol}_{g}\right).$ (215)
Lemma 6.1.
The operator $T$ is injective.
Proof.
For $f\in\Pi(M,\mathrm{vol}_{g})$ such that $f\neq 0$, let $v\in H$ be the solution of
$\begin{cases}-{\Delta}_{g}v=f&\text{in }M,\\ \partial_{\nu}v=0&\text{in }\partial M.\end{cases}$ (216)
The solution to this problem exists since $\int_{M}f=0=\int_{\partial M}\partial_{\nu}v$. In fact by regularity theory we can conclude that $v\in H$, and so we conclude that
$T(f)(v)=-\int_{M}f{\Delta}_{g}v\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\int_{M}f^{2}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}\neq 0.$ (217)
Thus $T(f)\neq 0$, concluding that $T$ is injective. ∎
Let us now prove the existence of $L^{2}$-weak limit eigenfunctions.
Proposition 6.2.
Let $(r_{n})$ be a sequence of positive numbers such that $r_{n}\to 0$. For any $n$ let $(\lambda_{k,{r_{n}}})$ be the eigenvalues of the operator $\tilde{\Delta}_{r_{n}}$ and let $(f_{k,r_{n}})$ be corresponding eigenfunctions. Then for any $k$ there exists a Laplace (resp. Neumann) eigenfunction $f$ of $(M,g)$ with associated eigenvalue $\mu$ such that, up to extracting subsequences, satisfy
$f_{k,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\rightharpoonup}}f,$ (218)
$\lambda_{k,r_{n}}\to C_{m}\,\mu,$ (219)
$\sup_{n}E_{r_{n}}(f_{k,r_{n}}-f)<+\infty.$ (220)
Proof.
By the proof of Theorem 1, in particular (54), there exists $\lambda\geq 0$ and a subsequence such that $\lambda_{k}(-\tilde{\Delta}_{r_{n}})\rightarrow\lambda$ up to subsequence. Since $\|f_{k,r_{n}}\|_{L^{2}(M)}=1$ for any $n$, there exists $f\in L^{2}(M,\mathrm{vol}_{g})$ such that the weak convergence (218) holds up to subsequence. Therefore, by Theorem 2, we get that for any $\psi\in C^{\infty}(M)$ (resp. $C^{\infty}_{\partial_{\nu}}(M)$),
$\int_{M}f\Delta_{g}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}$ $=\lim_{n}\frac{1}{C_{m}}\int_{M}f_{k,r_{n}}\tilde{\Delta}_{r_{n}}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\frac{1}{C_{m}}\lim_{n}\int_{M}\left(\tilde{\Delta}_{r_{n}}f_{k,r_{n}}\right)\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}$
$=-\frac{1}{C_{m}}\lim_{n}\lambda_{k}(-\tilde{\Delta}_{r_{n}})\int_{M}f_{k,r_{n}}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=-\frac{1}{C_{m}}\lambda\int_{M}f\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}.$
Moreover, since for any $n$ it holds that $\lambda_{k,r_{n}}>0$ and
$0=\int_{M}-\tilde{\Delta}_{r_{n}}f_{k,r_{n}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\int_{M}\tilde{\lambda}_{k,r_{n}}f_{k,r_{n}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}$
we get that $\int_{M}f_{k,r_{n}}=0$. Thus $\int_{M}f\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=0$ by weak convergence.
Now let $v\in W^{2,2}(M)$ be the solution of
$\begin{cases}-{\Delta}_{g}v=\frac{\lambda}{C_{m}}f&\text{ in }M\\ \partial_{\nu}v=0&\text{ in }\partial M\end{cases}$ (221)
satisfying $\int_{M}v=0$. Then we have that for $\psi\in H\cap C^{\infty}(M)$ then
$\int_{M}v{\Delta}_{g}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=-\frac{1}{C_{m}}\lambda\int_{M}f\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\int_{M}f{\Delta}_{g}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}.$ (222)
Since this is a dense subspace of $H$ and the functionals are continuous with respect to $W^{2,2}(M)$ in $\psi$, the equality holds for all $H$. Thus by Lemma 6.1 we conclude that $v=f$, and so $f$ satisfies
$\begin{cases}-{\Delta}_{g}f=\frac{\lambda}{C_{m}}f&\text{ in }M\\ \partial_{\nu}f=0&\text{ in }\partial M\end{cases}$ (223)
thus $f$ is a Neumann eigenfunction, and so it must be $C^{\infty}(M)$. Also since both $f_{k,r_{n}},f\in\Pi(M,\mathrm{vol}_{g})$, we know by Proposition 3.14 using triangle inequality of the inner product,
$E_{r_{n}}(f_{k,r_{n}}-f)^{\frac{1}{2}}\leq E_{r_{n}}(f_{k,r_{n}})^{\frac{1}{2}}+E_{r_{n}}(f)^{\frac{1}{2}}.$ (224)
We know that $E_{r_{n}}(f_{k,r_{n}})=\lambda_{k,r_{n}}\|f_{k,r_{n}}\|_{L^{2}(M)}=\lambda_{k,r_{n}}$ which is uniformly bounded. Also since $f\in C^{\infty}(M)$, by Lemma 5.1, we know that $E_{r_{n}}(f)$ is also uniformly bounded, concluding the proof. ∎
6.2. Energy comparison
Let us now compare the energy of a map defined on $M$ with the energy of the image of the map through a local chart parametrizing a neighborhood of an open subset of $\partial M$. To this aim, up to scaling, we consider a map $\Phi:(-1,1)^{m-1}\times[0,1)\to M$ which is a bi-Lipschitz homeomorphism onto its image. We set
$\mathcal{Q}\coloneqq(-1/2,1/2)^{m-1}\times[0,1/2).$ (225)
Lemma 6.3.
There exist constants $\tilde{C}=\tilde{C}(\Phi)>0$ and $\tilde{c}=\tilde{c}(\Phi)>0$ such that for any $f\in L^{2}(M)$, for any $r\in(0,1/2)$,
$\tilde{E}_{\tilde{c}r,\mathfrak{Q}}(f\circ\Phi)\leq\tilde{C}E_{r,\mathfrak{M}}(f).$ (226)
where $\mathfrak{Q}\coloneqq(\mathcal{Q},d_{\infty},\mathcal{L}^{m})$ and $\mathfrak{M}\coloneqq(M,d,\mu)$.
Proof.
We start by pointing out that there exist constants $c=c(\Phi)>0$ and $C=C(\Phi)>0$ such that for all $x\in\mathcal{Q}$ and $r\in(0,1/2)$,
$\Phi(Q_{cr}(x))\subset B_{r}(\Phi(x))\subset\Phi(Q_{Cr}(x)),$
$V(\Phi(x),r)\leq C\mathcal{L}^{m}(Q_{cr}(x)\cap\mathcal{Q}),$
$\det(g_{x})\geq 0,$
where $g_{x}$ is the metric in the coordinates given by $\Phi$. Then for any $x,y\in\mathcal{Q}$,
$\tilde{a}_{r,\mathfrak{M}}(\Phi(x),\Phi(y))$ $=1_{B_{r}(\Phi(x))}(\Phi(y))\left(\frac{1}{V(\Phi(x),r)}+\frac{1}{V(\Phi(y),r)}\right)$
$\geq 1_{Q_{cr}(x)}(y)\left(\frac{1}{C\mathcal{L}^{m}(Q_{cr}(x)\cap\mathcal{Q})}+\frac{1}{C\mathcal{L}^{m}(Q_{cr}(y)\cap\mathcal{Q})}\right)$
$=\frac{\tilde{a}_{cr,\mathfrak{Q}}(x,y)}{C}\,\cdot$
Thus
$\tilde{E}_{r,\mathfrak{M}}(f)$ $\geq\iint_{\Phi(\mathcal{Q})^{2}}\tilde{a}_{r,\mathfrak{M}}(p,q)\left(\frac{f(p)-f(q)}{r}\right)^{2}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(q)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(p)$
$=\iint_{\mathcal{Q}^{2}}\tilde{a}_{r,\mathfrak{M}}(\Phi(x),\Phi(y))\left(\frac{f(\Phi(x))-f(\Phi(y))}{r}\right)^{2}$
$\qquad\qquad\qquad\qquad\qquad\qquad\times\sqrt{\det(g_{x})\det(g_{y})}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
$\geq\iint_{\mathcal{Q}^{2}}\frac{c}{C}\tilde{a}_{cr,\mathfrak{Q}}(x,y)\left(\frac{f(\Phi(x))-f(\Phi(y))}{r}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
$=\frac{c}{C}E_{cr,\mathfrak{Q}}(f\circ\Phi).$
Taking $\tilde{c}=c$ and $\tilde{C}=\frac{C}{c}$, we obtain the result. ∎
6.3. Proof of Theorem 3
We are now in a position to prove Theorem 3. Recall the context of this result : $(r_{n})\subset(0,+\infty)$ is a sequence such that $r_{n}\to 0$, $(M^{m},g)$ is a compact, connected, smooth Riemannian manifold with $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$ ), $k$ is a positive integer, $\mu_{k}$ is the $k$-th lowest Laplace (resp. Neumann) eigenvalue of $\Delta_{g}$, and $f_{k,r_{n}}$ is an eigenfunction of $\tilde{\Delta}_{r_{n}}$ associated with the $k$-th eigenvalue $\lambda_{k}(-\tilde{\Delta}_{r_{n}})$ of this operator.
Proof.
Step 1. First we show strong $L^{2}$-convergence of the sequence $(f_{k,r_{n}})$. We proceed by contradiction. By Proposition 6.2, we can assume that there exist $\alpha>0$, $f\in L^{2}(M,\mu)$ which is a Neumann eigenfunction, and $(r_{n})\subset(0,+\infty)$ such that $r_{n}\to 0$ and
$f_{k,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\rightharpoonup}}f,\quad\quad\|f_{k,r_{n}}-f\|_{L^{2}(M)}^{2}\geq\alpha.$ (227)
Since $M$ is a compact manifold with boundary, up to scaling there exist finitely many bi-Lipschitz homeomorphisms $\{\Phi_{j}:(-1,1)^{m-1}\times[0,1)\rightarrow M\}_{j\in\{1,\ldots\ell\}}$ such that
$\bigcup_{j}\Phi_{j}(\mathcal{Q})=M,$ (228)
where $\mathcal{Q}$ is as in (225), and
$\bigcup_{j}\Phi_{j}\Big{(}(-1/2,1/2)^{m-1}\times\{0\}\Big{)}=\partial M.$ (229)
Then there exists $j\in\{1,\ldots,\ell\}$ such that, up to a subsequence,
$\inf_{n}\int_{\Phi_{j}(\mathcal{Q})}|f_{k,r_{n}}-f|^{2}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}\geq\frac{\alpha}{\ell}>0.$ (230)
From this we conclude that there exists $\tilde{\alpha}>0$ such that
$\inf_{n}\int_{\mathcal{Q}}|f_{k,r_{n}}-f|^{2}\circ\Phi_{j}\mathop{}\!\mathrm{d}\mathcal{L}^{m}\geq\tilde{\alpha}>0.$ (231)
Let us set $\Phi:=\Phi_{j}$. Then there exist $C,\tilde{C},\tilde{c}>0$ such that for any $n$,
[…]
By the weak convergence we also have
$h_{n}\coloneqq(f_{k,r_{n}}-f)\circ\Phi\stackrel{{\scriptstyle L^{2}}}{{\rightharpoonup}}0.$ (232)
Let us set $\overline{r}_{n}=\tilde{c}r_{n}$. For an integer $N$ to be chosen later, consider a decomposition of $\mathcal{Q}$ into $L_{N}$ disjoint subcubes $\{\tilde{Q}_{i}\}$ of size $1/N$. For any $x,y\in\mathcal{Q}$, we set
$a_{r}(x,y)\coloneqq\chi_{Q_{r}(x)\cap\mathcal{Q}}(y)\left(\frac{1}{\mathcal{L}^{m}(Q_{r}(x)\cap\mathcal{Q})}+\frac{1}{\mathcal{L}^{m}(Q_{r}(y)\cap\mathcal{Q})}\right),$ (233)
$a_{r,i}(x,y)\coloneqq\chi_{Q_{r}(x)\cap\mathcal{Q}_{i}}(y)\left(\frac{1}{\mathcal{L}^{m}(Q_{r}(x)\cap\mathcal{Q}_{i})}+\frac{1}{\mathcal{L}^{m}(Q_{r}(y)\cap\mathcal{Q}_{i})}\right),$ (234)
and we point out that for $x,y\in\mathcal{Q}_{i}$
$a_{r}(x,y)\geq\frac{1}{2^{m}}a_{r,i}(x,y).$ (235)
We also set for any $n$,
$\epsilon_{i,n,N}\coloneqq\int_{\mathcal{Q}_{i}}h_{n}\mathop{}\!\mathrm{d}\mathcal{L}^{m},\qquad\delta_{n,N}\coloneqq\max_{i}|\epsilon_{i,n,N}|.$ (236)
We obtain that for any $n$,
$\tilde{E}_{\overline{r}_{n},\mathfrak{Q}}(h_{n})$ $=\int_{\mathcal{Q}}\left(\int_{\mathcal{Q}}a_{\overline{r}_{n}}(x,y)\frac{\left(h_{n}(x)-h_{n}(y)\right)^{2}}{\overline{r}_{n}^{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (237)
$=\sum_{i}\int_{\mathcal{Q}_{i}}\left(\int_{\mathcal{Q}}a_{\overline{r}_{n}}(x,y)\frac{\left(h_{n}(x)-h_{n}(y)\right)^{2}}{\overline{r}_{n}^{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (238)
$\geq\frac{1}{2^{m}}\sum_{i}\int_{\mathcal{Q}_{i}}\left(\int_{\mathcal{Q}_{i}}a_{\overline{r}_{n},i}(x,y)\frac{\left(h_{n}(x)-h_{n}(y)\right)^{2}}{\overline{r}_{n}^{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (239)
$=\frac{1}{2^{m}}\sum_{i}\tilde{E}_{\overline{r}_{n},\mathcal{Q}_{i}}(h_{n})$ (240)
$=\frac{1}{2^{m}}\sum_{i}\tilde{E}_{\overline{r}_{n},\mathcal{Q}_{i}}(h_{n}-\epsilon_{i,n,N})$ (241)
$\geq\frac{1}{2^{m}}\sum_{i}\|h_{n}-\epsilon_{i,n,N}\|_{L^{2}(\mathcal{Q}_{i})}^{2}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathcal{Q}_{i}})$ (242)
$\geq\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\sum_{i}\|h_{n}-\epsilon_{i,n,N}\|_{L^{2}(\mathcal{Q}_{i})}^{2}$ (243)
$=\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\sum_{i}\left(\|h_{n}\|_{L^{2}(\mathcal{Q}_{i})}^{2}-2\epsilon_{i,n,N}\int_{\mathcal{Q}_{i}}h_{n}\mathop{}\!\mathrm{d}\mathcal{L}^{m}+\mathcal{L}^{m}(\mathcal{Q}_{i})\epsilon_{i,n,N}^{2}\right)$ (244)
$\geq\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\left(\|h_{n}\|^{2}_{L^{2}(\mathcal{Q})}-3L_{n}\delta_{n,N}\right)$ (245)
$\geq\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\left(\tilde{\alpha}-3L_{n}\delta_{n,N}\right).$ (246)
By Lemma 4.3, we choose $N$ big enough to ensure that for any $n$,
$\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}(1/N)})>C\tilde{C}\frac{2^{m+2}}{\tilde{\alpha}}\,\cdot$ (247)
By the weak convergence (232) we know that $\delta_{n,N}\rightarrow 0$, and so we can choose $n$ big enough to guarantee
$\delta_{n,N}<\frac{\tilde{\alpha}}{6L_{N}}\,\cdot$ (248)
With these choices we eventually get
$\tilde{E}_{r_{n}}(f_{k,r_{n}}-f)>C,$ (249)
which is a contradiction.
Step 2. Now we show that $\tilde{\lambda}_{k,r_{n}}\rightarrow\mu_{k}$, where $\mu_{k}$ is the $k$-th Neumann eigenvalue. Let $r_{n}\rightarrow 0$. We know by Proposition 6.2 that there exist eigenfunctions $f_{0},...,f_{k}$ with Neumann eigenvalue $\lambda_{0},...,\lambda_{k}$ such that
$f_{i,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\rightarrow}}f_{i},\quad\quad\forall i\in\{0,...,k\},$ (250)
$\tilde{\lambda}_{k,r_{n}}\rightarrow C_{m}\lambda_{k},$ (251)
and
$\lambda_{i}\leq\lambda_{k}\quad\quad\forall i\in\{0,...,k\}.$ (252)
Since $\langle f_{i,r_{n}},f_{j,r_{n}}\rangle=\delta_{i,j}$, we also have by strong convergence that $\langle f_{i},f_{j}\rangle=\delta_{i,j}$. Thus we have that
$V_{k+1}\coloneqq\mathrm{Span}(f_{0},\ldots,f_{k})\in\mathcal{G}_{k+1}(L^{2}(M,\mathrm{vol}_{g})),$ (253)
and so by equation (252), we conclude
$C_{m}\mu_{k}\leq\max_{f\in V_{k+1}}\frac{\langle\nabla f,\nabla f\rangle}{\|f\|_{L^{2}}}=C_{m}\lambda_{k}=\lim_{n}\tilde{\lambda}_{k,r_{n}}.$ (254)
This shows that $\liminf_{r\rightarrow 0}\tilde{\lambda}_{k,r_{n}}\geq C_{m}\mu_{k}$
To prove $\limsup_{r\rightarrow 0}\tilde{\lambda}_{k,r}\leq C_{m}\mu_{k}$, let $\{f_{0},\ldots,f_{k}\}$ be an $\langle\cdot,\cdot\rangle_{2}$-orthonormal family of Laplace (resp. Neumann) eigenfunctions associated with the eigenvalues $\{\mu_{0},\ldots,\mu_{k}\}$ respectively satisfying $\mu_{0}\leq...\leq\mu_{k}$. By elliptic regularity, we know that these functions belong to $C^{\infty}(M)$. Then Proposition 5.1 implies that given $\epsilon>0$, there exists $r_{\epsilon}>0$ such that for $r\in(0,r_{\epsilon})$,
$\left|\langle-\tilde{\Delta}_{r}f_{i},f_{j}\rangle_{2}-\delta_{i,j}C_{m}\mu_{j}\right|<\epsilon$ (255)
where $\delta_{i,j}$ is the usual Kronecker delta. Set $U\coloneqq\mathrm{Span}\left(f_{0},...,f_{k}\right)$ and
$v\coloneqq\sum_{i=1}^{k}a_{i}\psi_{i}$ (256)
for some $a=(a_{1},\ldots,a_{k})\in\mathbb{S}^{k-1}$. Then
$\left|\langle-\tilde{\Delta}_{r}v,v\rangle-\sum_{i=1}^{k}a_{i}^{2}C_{m}\mu_{i}\right|=\left|\sum_{i,j=1}^{k}a_{i}a_{j}\langle-\tilde{\Delta}_{r}f_{i},f_{j}\rangle-\sum_{i=1}^{k}a_{i}^{2}C_{m}\mu_{i}\right|\leq k^{2}\epsilon.$ (257)
Since $U$ is a $k+1$-dimensional subspace, we conclude that
$\tilde{\lambda}_{k,r}\leq\max_{v\in U}\frac{\langle-\tilde{\Delta}_{r}v,v\rangle}{\|v\|_{2}^{2}}\leq\max_{a\in\mathbb{S}^{k}}\sum_{i=1}^{k}a_{i}^{2}\mu_{i}+k^{2}\epsilon\leq\mu_{k}+k^{2}\epsilon.$ (258)
Take the limit superior as $r\to 0$ and then let $\epsilon\to 0$ to obtain $\limsup_{r\rightarrow 0}\tilde{\lambda}_{k,r}\leq C_{m}\mu_{k}$. Combined with Corollary 3.13, the latter implies the existence of $r_{k}>0$ such that $\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})\big{)}\geq\mu_{k}+1\geq\tilde{\lambda}_{k,r}$ for any $r\in(0,r_{k})$, so that $\tilde{\lambda}_{k,r}$ indeed coincides with $\lambda_{k}(-\tilde{\Delta}_{r})$. ∎